### v2

# Data Management For Large Language Models: A Survey

Abstract
Data plays a fundamental role in the training of Large Language Models (LLMs). Effective data management, particularly in the formulation of a well-suited training dataset, holds significance for enhancing model performance and improving training efficiency during pretraining and supervised fine-tuning phases. Despite the considerable importance of data management, the current research community still falls short in providing a systematic analysis of the rationale behind management strategy selection, its consequential effects, methodologies for evaluating curated datasets, and the ongoing pursuit of improved strategies. Consequently, the exploration of data management has attracted more and more attention among the research community. This survey provides a comprehensive overview of current research in data management within both the pretraining and supervised fine-tuning stages of LLMs, covering various noteworthy aspects of data management strategy design: data quantity, data quality, domain/task composition, etc. Looking toward the future, we extrapolate existing challenges and outline promising directions for development in this field. Therefore, this survey serves as a guiding resource for practitioners aspiring to construct powerful LLMs through effective data management practices. The collection of the latest papers is available at https://github.com/ZigeW/data_management_LLM.
Data Management For Large Language Models: A Survey
[…]
1 Introduction
Large Language Models (LLMs) have shocked the natural language processing (NLP) community with their strong performance and emergent abilities OpenAI (2023); Touvron et al. (2023a); Wei et al. (2022). According to previous studies Kaplan et al. (2020); Hoffmann et al. (2022), LLMs’ achievements depend heavily on self-supervised pretraining over processed vast volumes of text data. Recent research Zhou et al. (2023a); Ouyang et al. (2022) further enhances LLMs’ instruction-following ability and performance on downstream tasks through Supervised Fine-Tuning (SFT) on deliberately curated instruction datasets.
Organizing a well-suited training dataset using collected data, which we define as data management, is vitally important and challenging in both the pretraining and SFT stages of LLMs. In the pretraining stage, constructing datasets with high-quality data is essential for efficient training Jain et al. (2020); Gupta et al. (2021). To equip LLMs with diverse and comprehensive abilities, heterogeneous dataset composition with mixtures of domains is also required Gao et al. (2020); Longpre et al. (2023b); Shen et al. (2023). However, many prominent LLMs do not enclose Anil et al. (2023); OpenAI (2023) or only document the techniques used in the construction of their pretraining dataset Brown et al. (2020); Workshop et al. (2022); Touvron et al. (2023a), leaving the reason and effect of certain data management strategy absent. In the SFT stage, the performance and instruction-following abilities of LLMs are largely evoked by carefully constructed instruction datasets Sanh et al. (2022); Ouyang et al. (2022). Although a handful of instruction datasets/benchmarks have been proposed with human crowd sourcing Wang et al. (2022); Köpf et al. (2023), self-instruct Wang et al. (2023c); Taori et al. (2023) or collection of existing datasets Si et al. (2023); Anand et al. (2023), practitioners still find it confusing about the effect of instruction datasets on the performance of fine-tuned LLMs, leading to difficulties in choosing proper data management strategies in LLM SFT practices.
{forest}
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=center, font=, rectangle, draw=hidden-draw, rounded corners, align=left, text centered, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=13em,font=,, where level=2text width=14em,font=,, where level=3text width=11em,font=,, [ Data Management, ver [ Pretraining (§2), ver [ Data Quantity (§2.1) [ Scaling Laws [ Kaplan et al. (2020), DeepMind Chinchilla Scaling Law Hoffmann et al. (2022), leaf, text width=33em ] ] [ Data Repetition [ Villalobos et al. (2022), Muennighoff et al. (2023), Hernandez et al. (2022),
Xue et al. (2023), D4 Tirumala et al. (2023), leaf, text width=33em ] ] ] [ Data Quality (§2.2) [ Deduplication [ Lee et al. (2021), Kandpal et al. (2022), Silcock et al. (2022),
SemDeDup Abbas et al. (2023), leaf, text width=33em ] ] [ Quality Filtering [ Gao (2021), Kreutzer et al. (2022), Gunasekar et al. (2023), Li et al. (2023b),
RefinedWeb Penedo et al. (2023), Marion et al. (2023), Longpre et al. (2023b),
MiniPileKaddour (2023), leaf, text width=33em ] ] [ Toxicity Filtering [ Luccioni and Viviano (2021), Xu et al. (2021), Welbl et al. (2021),
Longpre et al. (2023b), leaf, text width=33em ] ] [ Social Bias [ Dodge et al. (2021), Meade et al. (2022), Gururangan et al. (2022)
Feng et al. (2023), leaf, text width=33em ] ] [ Diversity & Age [ Lee et al. (2023a), D2 Pruning Maharana et al. (2023), Longpre et al. (2023b), leaf, text width=33em ] ] ] [ Domain Composition (§2.3) [ Longpre et al. (2023b), CodeGen2 Nijkamp et al. (2023),
SlimPajama-DC Shen et al. (2023), DSIR Xie et al. (2023b),
DoReMi Xie et al. (2023a), DoGE Fan et al. (2023), leaf, text width=33em ] ] [ Data Management Systems (§2.4) [ Data-Juicer Chen et al. (2023a), Oasis Zhou et al. (2023c), leaf, text width=33em ] ] ] [ Supervised Fine-Tuning (§3), ver [ Data Quantity (§3.1) [ Ji et al. (2023), LIMA Zhou et al. (2023a), Yuan et al. (2023),
Chen et al. (2023b), DMT Dong et al. (2023), Song et al. (2023), leaf, text width=33em ] ] [ Data Quality (§3.2) [ Instruction Quality [ INSTRUCTEVAL Chia et al. (2023), LIMA Zhou et al. (2023a),
Ding et al. (2023), Wang et al. (2023d), Li et al. (2023a),
Instruction Mining Cao et al. (2023), Self-Refine Madaan et al. (2023),
SELF Lu et al. (2023a), SelFee Ye et al. (2023), leaf, text width=33em ] ] [ Instruction Diversity [ UltraChat Ding et al. (2023), LIMA Zhou et al. (2023a),
Alpaca Taori et al. (2023), #InsTag Lu et al. (2023b),
Self-Instruct Wang et al. (2023c), Explore-Instruct Wan et al. (2023), leaf, text width=33em ] ] [ Instruction Complexity [ #InsTag Lu et al. (2023b), WizardLM Xu et al. (2023a),
WizardCoder Luo et al. (2023), Orca Mukherjee et al. (2023),
Tree-Instruct Zhao et al. (2023b), CELLO He et al. (2023),
FollowBench Jiang et al. (2023), leaf, text width=33em ] ] [ Prompt Design [ Mishra et al. (2022), Khashabi et al. (2022), Gonen et al. (2022),
Yin et al. (2023b), Kung and Peng (2023), UIT Liang et al. (2023),
Weber et al. (2023), Gudibande et al. (2023), Song et al. (2023), leaf, text width=33em ] ] ] [ Task Composition (§3.3) [ Wei et al. (2021), Wang et al. (2022), Sanh et al. (2022), Chung et al. (2022),
Flan 2022 Longpre et al. (2023a), ELM Jang et al. (2023), Chen et al. (2023b)
DMT Dong et al. (2023), OPT-IML Iyer et al. (2022), Tulu Wang et al. (2023b), leaf, text width=34em ] ] [ Data-Efficient Learning (§3.4) [ AlShikh et al. (2023), Attendu and Corbeil (2023), Ivison et al. (2023),
Instrucion Mining Cao et al. (2023), AlpaGasus Chen et al. (2023c),
OpenChat Wang et al. (2023a), DiverseEvol Wu et al. (2023),
Dynosaur Yin et al. (2023a), MAmmoTH Yue et al. (2023),
DMT Dong et al. (2023), LoBaSS Zhou et al. (2023b),
Xu et al. (2023b), Data-Juicer Chen et al. (2023a), leaf, text width=34em ] ] ] ]
Figure 1: Taxonomy of research in data management for pretraining and supervised fine-tuning of Large Language Models (LLM).
To address these challenges, it is necessary to conduct a systematic analysis on LLM data management, including the rationale behind management strategy selection and its consequential effect, the evaluation of curated training datasets, and the pursuit of improved strategies. Therefore, this survey aims to provide a comprehensive overview of current research in LLM data management, as shown in Figure 1. In Section 2, we focus on LLM pretraining data management, including the research on data quantity, data quality, domain composition, and data management systems. In Section 3, we discuss the data quantity, data quality, task composition, and data-efficient learning in the SFT stage of LLMs. In Section 4, looking into the future, we present the existing challenges and promising future directions in training data management for LLMs. Through this survey, we are devoted to offering a guiding resource to practitioners attempting to build powerful LLMs with effective and efficient data management practices.
2 Pretraining of LLM
Data management is found to be important in the pretraining stage of many prominent LLMs OpenAI (2023); Touvron et al. (2023a); Wei et al. (2022). Understanding the effects of these data management strategies is also crucial for building strong LLMs. However, most of the prominent LLMs do not report their data management procedures or only report the strategies they adopted without discussing the reason for choosing the specific strategies. Thus, some researchers try to disclose the working scheme of data management in the pretraining stage of LLMs. In this section, we first review the research studying pretraining dataset scaling law with/without data repetition. Then, data quality regarding deduplication, quality filtering, toxicity filtering, social bias, and data diversity and age are explored. After that, domain composition and domain re-weighting methods are discussed. Finally, two recently proposed data management systems are introduced to implement the pretraining data management pipelines.
2.1 Data Quantity
The amount of data required for efficient pretraining of LLMs is an ongoing research topic in NLP communities. Scaling laws are proposed to depict the relationship between model size and training dataset size. According to the proposed scaling laws Kaplan et al. (2020); Hoffmann et al. (2022), with model size continuously increasing, the demand for more training data will also increase consistently. Thus, the exhaustion of text data draws researchers’ attention to data repetition in the LLMs’ pretraining.
2.1.1 Scaling Laws
Before the popularization of LLMs, the relationship between training dataset size and the performance of Transformer-based language models Vaswani et al. (2017) had already attracted researchers’ attention. Kaplan et al. (2020) use Transformers and cross-entropy loss to study the empirical scaling laws for language model performance and find the model performance has a power-law relationship with training dataset size or model size, respectively, when not bottlenecked by each other and the training compute budget:
$L(D)=\big{(}\frac{D_{c}}{D}\big{)}^{\alpha_{D}};\alpha_{D}\sim 0.095,D_{c}\sim 5.4\times 10^{13}$ (1)
$L(N)=\big{(}\frac{N_{c}}{N}\big{)}^{\alpha_{N}};\alpha_{N}\sim 0.076,N_{c}\sim 8.8\times 10^{13}$ (2)
where $L$ is the test loss, $D$ is the number of training tokens, $N$ is the number of model parameters, $\alpha_{D}$ and $\alpha_{N}$ is the power-law component for the scaling of $D$ and $N$ respectively, and $D_{c}$ and $N_{c}$ is constant numbers measured in the number of training tokens and non-embedding parameters respectively whose precise numerical value depends on vocabulary size and tokenization and does not have fundamental meaning.
Combining Equation 1 and  2, they derive Equation 3 to depict the dependence between model size $N$ and dataset size $D$. Fitting Equation 3, they get $\alpha_{D}=0.103$, $\alpha_{N}=0.076$, $N_{c}=6.4\times 10^{13}$ and $D_{c}=1.8\times 10^{13}$ and conclude that model performance improves predictably as long as the model size and training dataset size are scaled up simultaneously but encounters overfitting if either of them is fixed while the other increases. Given fixed compute budget $C$, they analyze the optimal allocation of $D_{opt}\sim C^{0.27}$ and $N_{opt}\sim C^{0.73}$, showing that the model size should increase faster than the training dataset size.
$L(N,D)=\big{[}\big{(}\frac{N_{c}}{N}\big{)}^{\frac{\alpha_{N}}{\alpha_{D}}}+\frac{D_{c}}{D}\big{]}^{\alpha_{D}}$ (3)
Following the power-law relationship proposed by Kaplan et al. (2020), Hoffmann et al. (2022) conduct experiments on much larger language models and arrive at a new scaling law:
$L(N,D)=E+\frac{A}{N^{\alpha}}+\frac{B}{D^{\beta}}$ (4)
where $E=1.69$, $A=406.4$, $B=410.7$, $\alpha=0.34$ and $\beta=0.28$. They also analyze the optimal allocation of $N_{opt}\sim C^{0.46}$ and $D_{opt}\sim C^{0.54}$. Hence, they draw a different conclusion that model size and pretraining dataset size should scale at roughly the same rate with a larger computing budget.
2.1.2 Data Repetition
While Kaplan et al. (2020) and Hoffmann et al. (2022) focus on scaling law with unique data trained only for one epoch, Hernandez et al. (2022) address the issue about text overlap in the training dataset and study the scaling law with a small fraction of repeated data. They observe a strong double descent phenomenon Nakkiran et al. (2021) caused by repeated data, where a peak of test loss appears in the middle range of repetition frequency, i.e., the number of epochs trained on repeated data. They also show that repeated data can cause a divergence from power-scaling law Kaplan et al. (2020) on model sizes larger than 100M parameters.
According to the scaling law, more training data is required as the model size grows, raising concerns about the exhaustion of high-quality training data Villalobos et al. (2022); Hoffmann et al. (2022). One straightforward way to overcome this issue is to train on data repeatedly. However, repeated data subsets may lead to performance degradation as claimed by Hernandez et al. (2022). Motivated by this contradiction, several works study the consequence of pretraining on the whole datasets repeatedly for multiple epochs. Muennighoff et al. (2023) find that with constrained data and fixed compute budgets, repeatedly training on the whole dataset up to 4 epochs only causes trivial harm to test loss compared to training on unique new data. They also propose a scaling law on repeated training depicting the diminishing of returns with more repetition and larger model sizes. Xue et al. (2023) also observe a multi-epoch degradation in model performance and find that dataset size, model parameters, and training objectives are the key factors to this phenomenon. They further find that commonly used regularization techniques are not helpful in alleviating multi-epoch degradation, except for dropout. Instead of simply repeating over the whole dataset, Tirumala et al. (2023) show that repeatedly training on carefully selected data can outperform that on randomly selected new data, whilst repeatedly training on randomly selected data cannot, suggesting a feasible way of repeating on intelligently selected data.
2.2 Data Quality
High-quality data is crucial in the training of machine learning tasks according to previous studies Jain et al. (2020); Gupta et al. (2021). In the pretraining of LLMs, quality assurance techniques are also adopted and usually form a data management pipeline Rae et al. (2021); Nguyen et al. (2023); Tirumala et al. (2023), including deduplication, quality filtering, and toxicity filtering. Other aspects like social bias, data diversity, and data age are also influential topics in the research community.
2.2.1 Deduplication
Deduplication is widely used in data management procedures of many prominent LLMs and preprocessing of publicly available datasets Brown et al. (2020); Workshop et al. (2022); Touvron et al. (2023a); Raffel et al. (2020). Lee et al. (2021) use N-gram similarity with MinHash Broder (1997) to detect duplications in training datasets and find that deduplication is beneficial in memorization mitigation, train-test overlap avoidance, and training efficiency improvement while keeping model perplexity. Kandpal et al. (2022) also show that deduplication can considerably lower the success rate of privacy attacks aiming at model memorization.
Among practices of deduplication, N-gram-and-hashing is the most commonly adopted technique Lee et al. (2021); Borgeaud et al. (2022); Rae et al. (2021). Silcock et al. (2022) compare it with two neural approaches (a contrastively trained bi-encoder and a "re-rank" style approach combining a bi- and cross-encoder) and conclude that neural approaches can significantly outperform traditional N-gram-and-hashing methods. Abbas et al. (2023) propose SemDeDup to remove semantic duplicates that lie closely in the pre-trained model embedding space and apply clustering to reduce the searching computation.
2.2.2 Quality Filtering
Quality filtering is another key step in constructing a high-quality pretraining dataset, because public datasets like Common Crawl  [^1] https://commoncrawl.org/, a large text corpus contains raw web page data, metadata extracts, and text extracts. and multilingual datasets Kreutzer et al. (2022) usually contain low-quality data that hampers the training of LLMs. Existing works usually perform quality filtering using a classifier Brown et al. (2020); Gao et al. (2020); Du et al. (2022); Touvron et al. (2023a), hand-crafted heuristics Yang et al. (2019); Raffel et al. (2020); Nijkamp et al. (2022) or threshold filtering using criterion like perplexity Wenzek et al. (2020); Muennighoff et al. (2023). Kaddour (2023) construct a subset of the Pile Gao et al. (2020) called MiniPile by filtering out low-quality embedding clusters.
Quality filtering is usually proven to be beneficial in model performance improvement Longpre et al. (2023b), despite the reduction of training data quantity and variety. Light-weight language model phi-1 Gunasekar et al. (2023) and phi-1.5 Li et al. (2023b) with 1.3B parameters trained on carefully selected high-quality data and synthetically generated data show outstanding performance on coding and commonsense reasoning tasks. Recently, Microsoft published a new 2.7B language model phi-2 Javaheripi and Bubeck (2023) with better language understanding and logic reasoning abilities. Penedo et al. (2023) construct the RefinedWeb dataset consisting of properly filtered and deduplicated high-quality web data, outperforming models trained on the Pile Gao et al. (2020). However, Gao (2021) finds that aggressive filtering can lead to performance degradation on a wide range of tasks for GPT-like LLMs because the filtering proxy objectives are not representative enough of the true objective. To address this issue, Marion et al. (2023) comprehensively examines the combinations of three data quality estimators (perplexity, Error L2-Norm (EL2N), and memorization factor), three proportions of remaining data (top, middle, and bottom proportion ordered by corresponding estimators) and different percentage of remaining data through data pruning. Surprisingly, they find that pruning datasets based on perplexity and retaining the middle proportion of data performs far better than more complicated techniques like memorization. However, no combinations of pruning strategies achieve consistently high performance.
2.2.3 Toxicity Filtering
Toxicity refers to the text content which is "rude, disrespectful, or unreasonable language that is likely to make someone leave a discussion" Gehman et al. (2020); Welbl et al. (2021). As raw text corpora usually contain toxic text Luccioni and Viviano (2021); Longpre et al. (2023b), toxicity filtering aims to remove text with undesirable toxic text in the pretraining datasets, further preventing LLMs from generating toxic utterances. Similar to quality filtering, heuristic and rule-based filtering Lees et al. (2022); Gargee et al. (2022); Friedl (2023) and N-gram classifiers Raffel et al. (2020) are usually adopted as toxicity filters. Although effective in model detoxifying, Longpre et al. (2023b) discover that toxicity filtering reduces the risk of toxic generation by sacrificing model generalization and toxicity identification ability. Moreover, Xu et al. (2021) and Welbl et al. (2021) both find that training dataset detoxification leads to the marginalization of minority groups like dialects and minority identity mentions.
2.2.4 Social Bias
Besides the marginalization of minority groups caused by data detoxifying, several works Kurita et al. (2019); Nangia et al. (2020); Meade et al. (2022); Feng et al. (2023) find that pre-trained LLMs can capture social biases contained in the large amounts of training text. Evaluating on the C4.EN Raffel et al. (2020) dataset, Dodge et al. (2021) recommend documenting the social biases and representational harms as well as excluded voices and identities in large web text corpora. Using a dataset of U.S. high school newspaper articles, Gururangan et al. (2022) also argue that the quality filters used for GPT-3 Brown et al. (2020) prefer newspapers published by larger schools located in wealthier, educated, and urban ZIP codes, leading to a language ideology. Feng et al. (2023) conduct a comprehensive case study focusing on the effects of media political biases in the pretraining corpus on the fairness of hate speech detection and misinformation detection w.r.t. partisan leanings and how it is propagated to language models even further to downstream tasks.
2.2.5 Diversity & Age
There are also works focusing on other aspects of data management in the pretraining stage of LLMs. For example, Lee et al. (2023a) show that the format diversities of publicly available pretraining datasets are high when measured by the recently proposed Task2Vec diversity coefficient Miranda et al. (2022). They also demonstrate the effectiveness of the diversity coefficient by showing its alignment with the number of latent concepts in the text and suggest using it to build more diverse datasets. Maharana et al. (2023) proposes a novel pruning method D2 Pruning to balance data diversity and difficulty in data selection by representing a dataset as an undirected graph with samples as nodes and difficulty scores as node properties. Each sample is connected with its k nearest neighbors weighted by distances in the embedding space. Then, a forward and reverse message passing strategy is adopted to select a subgraph enveloping both diverse and difficult data samples.
Longpre et al. (2023b) explore the age of the evaluation dataset and draw conclusions that the temporal shift between evaluation and pretraining data will lead to inaccurate performance estimation and the temporal misalignment might not be overcome by fine-tuning, especially for larger models.
2.3 Domain Composition
Public available pretraining datasets usually contain mixtures of data collected from multiple sources and domains. For example, the Pile Gao et al. (2020) contains web documents from Common Crawl, Wikipedia, Books, and collections from medical, academic, coding and math, legal, and social resources. Many prominent models are also trained on a mixture of data from different domains. For example, LaMDA Thoppilan et al. (2022) is trained on dialogs data from public forums, C4 data, Wikipedia (English), web documents, and code documents from programming-related Q&A sites, tutorials, etc.
Efforts are made to explore the impact of domain mixtures on the pre-trained model performance. Longpre et al. (2023b) group the Pile Gao et al. (2020) data into nine domains and conduct ablate-one-at-a-time experiments to show the impact of different domains. They draw conclusions that the domains with high quality (Books) and high diversity (Web) are broadly helpful, and it is beneficial to include as many data sources as possible even though they are less relevant to the downstream tasks. SlimPajama-DC Shen et al. (2023) arrives at the same point that including all domains typically achieves better results than deliberately selected domain combinations, given that global deduplication is conducted to remove overlaps among different domain datasets. Both Longpre et al. (2023b) and Shen et al. (2023) agree that specific mixtures may excel in evaluation benchmarks for targeted tasks, but the former claim that the inclusion of diverse web domains may perform better than specific mixtures in certain tasks. CodeGen2 Nijkamp et al. (2023) studies the impact of mixtures of programming and natural languages on model performance and finds that models trained with mixtures do not perform better than but closely to domain-matched models given the same compute budget.
Several methods are also proposed to find the proper domain composition weights. DSIR Xie et al. (2023b) formulates the problem as a distribution matching problem given a large raw unlabeled dataset and some unlabeled target samples. Specifically, it leverages the classic importance resampling approach Rubin (1988) and estimates the importance weights using n-gram features and KL reduction. Without knowledge of downstream tasks or target distributions, DoReMi Xie et al. (2023a) trains a small proxy model using Group Domain Robust Optimization (Group DRO) Oren et al. (2019); Sagawa* et al. (2020) to generate domain weights. It improves model performance on all domains by up-weighting domains with the largest loss gap between the evaluated model and a pre-trained reference model. Improved from DoReMi Xie et al. (2023a), Fan et al. (2023) propose DoGE which reweights training domains to minimize the average validation loss across all training domains or on a specific unseen domain. The key component of their generalization objective is a gradient-based generalization estimation function measuring the contribution of each domain to other domains. Then, domains contributing higher to learning other domains will receive larger weights, as well as those difficult to learn.
2.4 Data Management Systems
Addressing the difficulty in pretraining data management, integrated data management systems are beneficial for LLM practitioners with different demands. Chen et al. (2023a) provides a data processing system Data-Juicer featuring the generation of diverse data recipes. They provide over 50 versatile data management operators and dedicated tools targeting zero-code data processing, low-code customization, and off-the-shelf data processing components. A timely feedback loop at multiple development stages of data recipes and LLMs is also supported. Zhou et al. (2023c) also propose a pretraining data curation and assessment system Oasis, which contains an interactive modular rule filter module, a debiased neural quality filter module, an adaptive document deduplication module, and a holistic data assessment module.
3 Supervised Fine-Tuning of LLM
Based on the general knowledge and capabilities learned in the pretraining stage, supervised fine-tuning (SFT) is proposed to further improve LLMs with instruction-following ability and alignment with human expectations Wei et al. (2021); Sanh et al. (2022); Ouyang et al. (2022). Many efforts have been made to construct instruction data using human crowd-sourcing Wang et al. (2022); Köpf et al. (2023), self-instruct Wang et al. (2023c); Taori et al. (2023) or collection of existing datasets Si et al. (2023); Anand et al. (2023). Although LLMs fined-tuned with existing instruction datasets have achieved remarkable performance in various NLP tasks, the impacts of instruction data management on fine-tuned model performance are still under debate. Consistent with previous discussion regarding LLM pretraining, in this section, we summarize the research explorations in LLM SFT into data quantity, data quality (instruction quality, diversity, complexity, prompt design), and task composition. Data-efficient SFT is also included to discuss current efforts on efficient SFT from the data aspect.
3.1 Data Quantity
The explorations of the relationship between scaling instruction data quantity and fine-tuned model performance diverge in two directions. One branch of research focuses on scaling down the instruction data quantity to improve training efficiency. For example, LIMA Zhou et al. (2023a) carefully curated 1,000 high-quality samples and experimentally justified their hypothesis that only limited instruction tuning data is needed to expose the knowledge and capabilities that the LLM has already acquired during pretraining. Chen et al. (2023b) observe that maybe a single instruction is sufficient for single task-specific LLM fine-tuning, and 16K samples with 1.9M tokens may be sufficient to train a model specialized in the natural language inference (NLI) task. Another branch of research argues that scaling up the instruction data quantity is crucial for success Wei et al. (2021); Sanh et al. (2022).
Addressing this conflict, several works attempt to analyze the scaling pattern for different tasks or model abilities. Ji et al. (2023) conduct an empirical study on 12 major real-world online user cases and show that scaling up the instruction data leads to continuous improvement in tasks such as extraction, classification, closed QA, and summarization while leading to little improvement in tasks such as math, code, and chain-of-thought. Disagree with Ji et al. (2023), Dong et al. (2023) find that general ability can be enhanced with about 1,000 samples and improves slowly after then, while mathematical reasoning and code generation improve consistently with the increasing of data amount. Similarly, Yuan et al. (2023) observe a log-linear relation between instruction data amount and models’ mathematical reasoning performance, but stronger pre-trained models improve less with larger fine-tuning datasets. Song et al. (2023) conduct experiments covering ten distinct in-domain abilities and three out-of-domain abilities and show that most abilities are consistent with data scaling. Still, each ability develops at different paces during instruction tuning, while some abilities show completely different patterns.
3.2 Data Quality
Data quality is always a focal point in the SFT of LLMs, addressing instruction quality, diversity, complexity, and prompt design. Here, we focus more on the management and analysis of existing instruction data instead of instruction generation methods discussed in previous surveys Zhang et al. (2023b); Wang et al. (2023e).
3.2.1 Instruction Quality
Many researchers have found that the quality of instruction data is one of the most important factors in improving model performance Chia et al. (2023); Zhou et al. (2023a); Ding et al. (2023). During the construction of instruction data, there is usually a filtering step to select high-quality instructions generated by models. Wang et al. (2023d) use perplexity as the criterion to select the most appropriate instructions from the pool of candidate instructions generated by open-source models. Cao et al. (2023) propose an automatic data selector InstructionMining to evaluate instruction data quality without human experts’ interventions. They first hypothesize that the inference loss of a fine-tuned model on an evaluation set can serve as a proxy for data quality. Then, they use a set of natural language indicators to predict the inference loss to save the efforts in actually fine-tuning an LLM, including the number of tokens in tokenized inputs and outputs, reward score Köpf et al. (2023), perplexity, Measure of Textual Lexical Diversity McCarthy and Jarvis (2010), distance to approximate i-th nearest neighbors Dong et al. (2011) in SentenceBERT Reimers and Gurevych (2019) embedding space, and the naturalness, coherence, and understandability scores provided by UniEval dialogue model Zhong et al. (2022).
Instead of using indicators to filter low-quality instructions, several works Li et al. (2023a); Lu et al. (2023a); Ye et al. (2023); Madaan et al. (2023) leverage the power of fine-tuned LLM itself to evaluate the quality of instructions. Li et al. (2023a) assign quality scores to augmented instructions using the language model and iteratively improve model prediction. Similarly, SELF Lu et al. (2023a) and Self-Refine Madaan et al. (2023) prompts LLM to provide self-feedback on their own responses. Strong LLMs like ChatGPT are also adopted as quality judges during the instruction collection process Ye et al. (2023).
3.2.2 Instruction Diversity
The intention and semantic diversity of instructions is another important factor that has shown a positive effect on model performance improvement Zhou et al. (2023a); Ding et al. (2023); Taori et al. (2023). Self-Instruct Wang et al. (2023c) adopts ROUGE-L similarity to filter out the newly generated instructions that are too similar to the existing ones. To better evaluate the instruction diversity of SFT datasets, #InsTag Lu et al. (2023b) is proposed as an open-set fine-grained tagger using ChatGPT  [^2] https://chatgpt.openai.com/. Specifically, it first prompts ChatGPT to provide tags for given queries in an open setting, then performs a normalization procedure to deal with the noise in the raw tagging, including low-frequency filtering, rule aggregation, semantic aggregation with embedding clustering, and association aggregation that merges associated tags together. With the generated tags, they quantify instruction diversity as the unique tag coverage rate for the overall tag set. They also analyze popular open-set SFT datasets and show that larger dataset size tends to be more diverse and induces higher performance.
Though important, diversity can be challenging in domain-specific tasks due to data constraints. Wan et al. (2023) propose an approach called Explore-Instruct to enlarge the data coverage through active exploration via LLMs. Explore-Instruct starts from representative domain user cases and searches the variations and possibilities by looking ahead into potential fine-grained sub-tasks and backtracking alternative branches in the search space.
3.2.3 Instruction Complexity
The complexity of instructions also attracts researchers’ attention, especially in developing LLMs with complex instruction-following and reasoning abilities Xu et al. (2023a); Luo et al. (2023); Mukherjee et al. (2023). Several works endeavor to quantify and evaluate instruction complexity. Using aforementioned tags, #InsTag Lu et al. (2023b) quantifies complexity as the average tag number assigned to queries in a dataset. He et al. (2023) evaluate complex instruction with eight features, i.e., multi-turn, length, noise, and heterogeneous information for input text, multi-tasking, semantics, formats, and quantity constraints for task description.
To delve into the exploration of instruction complexity, Zhao et al. (2023b) propose Tree-Instruct to enhance the complexity of instruction data controllably. It treats the instruction as a semantic tree and constructs new complex instructions by adding nodes to the tree. Thus, the complexity of instruction can be controlled by adjusting the number of added nodes. Through experiments, they find that increased complexity can lead to continuing performance improvement. What’s more, the improvement does not come from the increased number of tokens, as a few complex instructions still outperform diverse but simple instructions under the same token budget. Curriculum instruction tuning ranging from easy to difficult might not be as helpful as expected, indicating the necessity of enhancing complexity. Another method of increasing complexity is also proposed. Evol-Instruct Xu et al. (2023a); Luo et al. (2023) rewrite instructions step by step with operations such as increasing reasoning, adding constraints, in-breadth evolving, deepening, and complicating input with code and table. Similarly, Jiang et al. (2023) incrementally augment instructions with constraints on content, situation, style, format and example and propose FollowBench to evaluate LLMs’ constraint following ability.
3.2.4 Prompt Design
Current instructions are either heuristically designed by human Wang et al. (2022); Köpf et al. (2023) or synthetically generated by prominent models Peng et al. (2023); Ding et al. (2023). However, the same intention and semantic meaning can be phrased into various prompts, and the choice of prompts can cause significant model performance variation Gonen et al. (2022); Weber et al. (2023). Thus, what kinds of instruction prompts are better for LLM training might be vital yet have not been fully explored.
Early attempts include manual reformulation of prompts into the ones easier to follow for language models Mishra et al. (2022) and choosing prompts with the lowest perplexity to get the most significant gains in model performance Gonen et al. (2022). Khashabi et al. (2022) surprisingly find that the discretized interpretation of continuous prompts is not always consistent with the discrete prompts describing the same task as heuristically expected. To know which parts of the instructions are most important in LLM fine-tuning, Yin et al. (2023b) and Kung and Peng (2023) both conduct ablation analysis removing contents in task definitions. Yin et al. (2023b) find that removing the descriptions of task output, especially the label information, might be the only reason for performance degradation. They also propose an automatic task definition compression algorithm to remove almost half or more of the tokens while improving model performance. Kung and Peng (2023) also remove all semantic components in task definitions but the output space information. They achieve comparable model performance using the modified task definitions and delusive examples containing incorrect input-output mappings. Based on this finding, they cast doubts on the performance gain of fine-tuned models and state that the model may only learn superficial patterns during instruction tuning. Addressing the issue of instruction format inconsistency, Liang et al. (2023) develop a format transfer framework UIT to transfer instructions from different datasets into unified formats automatically.
Besides the choice of phrasing, the generation source of prompts is another factor in prompt design. Gudibande et al. (2023) raise questions on fine-tuning a weaker language model on outputs of a stronger model and find that the imitation model might adapt to mimic the stronger model’s style but not its functionality, indicating the potential challenge of closing the performance gap between open- and close-sourced through imitation instruction tuning. Similarly, Song et al. (2023) also observe that human-designed data can outperform synthetically generated data from GPT-4 OpenAI (2023) to a relatively large extent.
3.3 Task Composition
Since LLMs have shown surprisingly emergent abilities in handling various NLP tasks, multitask fine-tuning appears to be a promising approach to improve LLMs’ generalization performance on unseen tasks. The benefits of increasing the number of tasks in SFT have been experimentally proven on models with different sizes ranging from 3B to 540B parameters Wang et al. (2022); Sanh et al. (2022); Wei et al. (2021); Chung et al. (2022).
Besides the scaling of the number of tasks, the mixture ratio of different instruction benchmarks and task balancing is also found to be critical for effective instruction fine-tuning Iyer et al. (2022); Longpre et al. (2023a). Dong et al. (2023) focus on task composition among mathematical reasoning, code generation, and general human-aligning abilities. Compared with individual source data, They find that model abilities are improved when the mixed data amount is small but decreased when the mixed data amount is large. The results indicate that conflicts exist among abilities with larger amounts of mixed data. To further explain the conflicts, they vary the ratio of general and specialized data and conclude that the impact of data ratio might lie in the similarity degree of data format and data distribution among different SFT tasks.
Different from compositing multiple tasks together, some works claim that LLM tuned on a single task data can outperform LLM tuned on multiple tasks Jang et al. (2023); Chen et al. (2023b). Jang et al. (2023) state that training expert LLMs to form an expert library can outperform multitask language models by avoiding negative task transfer, continually learning new tasks without catastrophic forgetting, and improving compositional abilities. Wang et al. (2023b) conduct analysis on factual knowledge, reasoning, multilinguality, coding, and open-ended instruction following abilities of models trained with 12 instruction datasets and experimentally show that different instruction datasets may correspond to specific abilities. Also, it seems that winning across all evaluations using a single dataset or combination is challenging.
3.4 Date-Efficient Learning
Based on explorations of the impact of data quantity, data quality, and task composition on model performance discussed previously, some works propose to fine-tune LLM more efficiently with subset selection or specially designed fine-tuning strategy addressing different aspects of instruction data.
Data Quantity
AlShikh et al. (2023) introduce a metric to measure LLMs’ instruction-following ability. A binary classifier is used to predict whether the response is "answer-like" or not. Then, Instruction Following Score (IFS) is defined as the percentage of "answer-like" responses in the responses acquired by the whole instruction dataset. They suggest using this metric as an early stopping criterion without tuning on the full-sized dataset. Based on observations of different scaling patterns for different abilities, Dong et al. (2023) propose Dual-stage Mixed Fine-tuning (DMT) strategy to learn specialized abilities and general abilities sequentially while keeping a small proportion of specialized data to prevent forgetting.
Data Quality
Several works focus on selecting a subset of instruction data with the highest quality. Cao et al. (2023) adopt BlendSearch Wang et al. (2020) to automatically select the best subset. AlpaGasus Chen et al. (2023c) uses strong LLMs as auto-graders and selects data with scores above a threshold in the Alpaca dataset Taori et al. (2023). Motivated by the computational overhead caused by data pruning prior to fine-tuning, Attendu and Corbeil (2023) propose a dynamic data pruning method that periodically filters out unimportant examples during SFT using extended versions of EL2N metric Paul et al. (2021); Fayyaz et al. (2022). Without discarding data samples, OpenChat Wang et al. (2023a) considers the general SFT data as a mixture of a small amount of expert data and a large amount of sub-optimal data without any preference labels. Then, Conditioned-RLFT strategy is proposed, which treats different data sources as coarse-grained reward labels and optimizes the LLM as a class-conditioned policy.
To enhance instruction diversity in the chosen subsets, DiverseEvol Wu et al. (2023) uses an iterative data sampling technique that selects new data points with maximized distances from any existing ones in model embedding space.
Task Composition
Given a small amount of target task data, Ivison et al. (2023) finds the relevant multitask subsets for fine-tuning according to the similarity between the pre-trained model’s representations. Similarly, Dynosaur Yin et al. (2023a) treats task selection based on instruction representations as a replay strategy in continual learning scenarios to mitigate catastrophic forgetting issues and improve generalization to unseen tasks.  Yue et al. (2023) builds math generalist models MAmmoTH through instruction tuning on a unique hybrid of chain-of-thought and program-of-though rationales in math.
Others
Zhou et al. (2023b) introduce learnability as a new dimension of SFT data selection that data can be learned more effectively by the model are preferable and data lacking informative content or excessively demanding for the model should be avoided. They also propose LoBaSS to select SFT data using learnability as the principal criterion measured by the loss difference between fine-tuned and pre-trained models. Xu et al. (2023b) propose a contrastive post-training technique via contrastive pairs from LLMs with different levels of abilities. They also use a data curriculum scheme where the model learns progressively from the "easier part" to the "harder part". Data-Juicer Chen et al. (2023a) implements pipelines for LLM fine-tuning and operators for users to compose different data management recipes, as well as the evaluation of consequence model performance.
4 Challenges and Future Directions
The exploration of data management and its impact on LLM pretraining and SFT is still an ongoing task. In this section, we point out several challenges and corresponding future directions in training data management studies for LLMs.
Comprehensive and Fine-grained Understanding
As discussed in previous sections, many efforts have been made to understand the impacts of data management on different training stages addressing different aspects. While current studies contribute valuable pieces to the puzzle, a comprehensive understanding of the entire picture is still lacking. Moreover, explorations using different datasets and models on different tasks may lead to contradictory conclusions, e.g., the trade-off between quality and toxicity filtering Longpre et al. (2023b), detoxification v.s. debiasing Xu et al. (2021); Welbl et al. (2021), fine-tuning with a few high-quality data Zhou et al. (2023a); Chen et al. (2023b) v.s. data scaling Wei et al. (2021); Sanh et al. (2022), task composition Wang et al. (2022); Chung et al. (2022) v.s. expert models Jang et al. (2023); Wang et al. (2023b), etc. Hence, more fine-grained understanding is required to solve these conflicts.
General Data Management Framework
Although Data-Juicer Chen et al. (2023a) and Oasis Zhou et al. (2023c) propose data management systems to compose various data recipes in either the pretraining or SFT stage of LLM, practitioners still need to spend efforts on finding suitable datasets. Constructing a general data management framework suitable for a broad range of applications to reduce data management costs is an urgent and worthy future direction in developing and promoting LLMs.
Data Efficiency Through Data Management
Training more powerful LLMs with less and more effective training data is always an ongoing pursuit in the training of LLMs. As discussed in Section 3.4, works have been proposed to fine-tune LLMs to achieve various abilities in a data-efficient way. Further exploration in data management is expected to offer good opportunities for achieving better data efficiency.
Data Curriculum
Besides choosing better training data, data curriculum addressing the arrangement of data learning orders is also an important part of data management, e.g., learning from general abilities to target abilities or from easier tasks to harder tasks. There are a few works focusing on data curriculum in the training of LLMs Xu et al. (2023b); Dong et al. (2023); Yin et al. (2023a). Although effective in practice, there is still a lack of analysis of data curriculum strategies.
Conflict Data Separation
In the collection of training data, conflicts among the responses to the same query may exist. For example, given the same query, LLMs playing different roles may generate different responses. Mixing these samples together could lead to negative impacts on model performance because of the response conflicts. Thus, how to separate and effectively learn from these data samples is an interesting topic in the future.
Multimodal Data Management
Current research in data management mostly focuses on natural language processing. With the application of LLMs extending to multimodalities like vision, audio, and so on, the construction of multimodality datasets becomes more and more important. The proposed multi-modal LLMs usually construct their own instruction-tuning datasets collected from benchmark adaptation Zhang et al. (2023a); Gao et al. (2023) or self-instruction Pi et al. (2023); Yang et al. (2023b). The hybrid composition of language-only and multimodal data is also adopted in some works Dai et al. (2023); Zhao et al. (2023c). It is interesting to see the impacts of multimodal data management on the performance of fine-tuned multimodal LLMs, such as the data scaling law in multimodal instruction fine-tuning, the quality-control techniques in multimodal dataset construction, task balancing in multitask multimodal training, and so on.
Hallucinations
Despite their strong power, LLMs are notorious for their hallucinations, i.e. the generation of input-, context- or fact-conflicting contents Zhang et al. (2023c). Several works in hallucination trace down the occurrence of hallucination to the lack of pertinent knowledge and the internalization of false knowledge from the pretraining corpora Li et al. (2022); McKenna et al. (2023); Dziri et al. (2022). To mitigate hallucination, the curation of pretraining corpora is adopted by many LLMs, mainly focusing on the extracting of high-quality data, e.g., GPT-3 Brown et al. (2020), Llama 2 Touvron et al. (2023b), and Falcon Penedo et al. (2023). The manually curated Zhou et al. (2023a) and automatically selected Chen et al. (2023c); Cao et al. (2023); Lee et al. (2023b) high-quality instruction data are also experimentally shown to be effective in reducing hallucination during the SFT stage. It can be seen from the previous research that data management in both the pretraining and SFT stages can be a promising solution to hallucination.
Social Biases and Fairness
The problem of social biases and unfairness existing in current pretraining datasets and their impacts on pre-trained LLMs have been addressed in several works as discussed in Section 2.2.4. However, there is still a large gap between current prominent LLMs and ideal LLMs without social biases. Many questions are worth exploring, such as how to mitigate the potential biases in pretraining datasets, the existence of bias in the SFT datasets, and whether it is feasible to reduce social bias through SFT.
5 Related Surveys
As LLMs draw more and more attention, a handful of surveys have been published or preprinted addressing different aspects of the life circle of LLMs. Related to our work, several of them also include parts of the data preparation process in the pretraining or SFT of LLM. Zhao et al. (2023a) review the development of LLMs and the latest advancements covering a wide range of topics. Yang et al. (2023a) also provide an overview of the LLM evolution and discuss the related techniques from model, data, and downstream tasks. Also concentrating on data, Zha et al. (2023) introduce data-centric AI and its related tasks and methods for general machine learning models instead of LLMs. Zhang et al. (2023b) survey the instruction tuning of LLMs and its related methodologies, data construction, applications, and so on. Wang et al. (2023e) review the technologies aligning LLMs with human expectations including data collection, training methodologies, and model evaluation.
Different from previous surveys, this survey provides a systematic and detailed overview of data management at both the pretraining and SFT stages of LLMs. We focus on the proper organization of training datasets and the effects of different data management strategies, providing a guiding resource for practitioners aiming to build powerful LLMs through effective and efficient data management.
6 Conclusions
This paper overviews data management in the training of LLMs. We discuss the pretraining and supervised fine-tuning stage of LLM successively and summarize the up-to-date research efforts into data quantity, data quality, and domain/task composition for each stage. Data management systems in the pretraining stage and Data-efficient learning in the supervised fine-tuning stage are also discussed. Finally, we highlight several challenges and promising future directions for LLM training data management. We hope this survey can provide insightful guidance for practitioners and inspire further research in effective and efficient data management for the development of LLMs.

### v3

@@
-Data Management For Large Language Models: A Survey
+Data Management For Training Large Language Models: A Survey
 Abstract
-Data plays a fundamental role in the training of Large Language Models (LLMs).
-Effective data management, particularly in the formulation of a well-suited training dataset, holds significance for enhancing model performance and improving training efficiency during pretraining and supervised fine-tuning phases.
-Despite the considerable importance of data management, the current research community still falls short in providing a systematic analysis of the rationale behind management strategy selection, its consequential effects, methodologies for evaluating curated datasets, and the ongoing pursuit of improved strategies.
+Data plays a fundamental role in training Large Language Models (LLMs).
+Efficient data management, particularly in formulating a well-suited training dataset, is significant for enhancing model performance and improving training efficiency during pretraining and supervised fine-tuning stages.
+Despite the considerable importance of data management, the underlying mechanism of current prominent practices are still unknown.
 Consequently, the exploration of data management has attracted more and more attention among the research community.
-This survey provides a comprehensive overview of current research in data management within both the pretraining and supervised fine-tuning stages of LLMs, covering various noteworthy aspects of data management strategy design: data quantity, data quality, domain/task composition, etc.
-Looking toward the future, we extrapolate existing challenges and outline promising directions for development in this field.
-Therefore, this survey serves as a guiding resource for practitioners aspiring to construct powerful LLMs through effective data management practices.
-The collection of the latest papers is available at https://github.com/ZigeW/data_management_LLM.
-Data Management For Large Language Models: A Survey
-[…]
+This survey aims to provide a comprehensive overview of current research in data management within both the pretraining and supervised fine-tuning stages of LLMs, covering various aspects of data management strategy design.
+Looking into the future, we extrapolate existing challenges and outline promising directions for development in this field.
+Therefore, this survey serves as a guiding resource for practitioners aspiring to construct powerful LLMs through efficient data management practices.
+Data Management For Training Large Language Models: A Survey
+Zige Wang1,2thanks:  Work done during Zige Wang’s internship at Huawei Noah’s Ark Lab.
+Wanjun Zhong2thanks:  Corresponding author (zhongwanjun1@huawei.com)  Yufei Wang2  Qi Zhu2  Fei Mi2  Baojun Wang2 Lifeng Shang2Xin Jiang2Qun Liu2 1School of Computer Science, Peking University 2Huawei Noah’s Ark Lab zigewang@stu.pku.edu.cn {zhongwanjun1, wangyufei44, zhuqi41, mifei2, puking.w}@huawei.com {Shang.Lifeng, Jiang.Xin, qun.liu}@huawei.com
 1 Introduction
@@
 (2022) further enhances LLMs’ instruction-following ability and performance on downstream tasks through Supervised Fine-Tuning (SFT) on deliberately curated instruction datasets.
-Organizing a well-suited training dataset using collected data, which we define as data management, is vitally important and challenging in both the pretraining and SFT stages of LLMs.
+To construct suitable training datasets, data management is vitally important and challenging in both the pretraining and SFT stages of LLMs, which we define as following:
+Data management: the process of organizing a well-suited training dataset with collected data, including the data selection, combination and utilization strategies, and the evaluation of the chosen strategies.
 In the pretraining stage, constructing datasets with high-quality data is essential for efficient training Jain et al.
@@
 However, many prominent LLMs do not enclose Anil et al.
-(2023); OpenAI (2023) or only document the techniques used in the construction of their pretraining dataset Brown et al.
-(2020); Workshop et al.
+(2023); OpenAI (2023) or only document Brown et al.
+(2020); Le Scao et al.
 (2023); Touvron et al.
-(2023a), leaving the reason and effect of certain data management strategy absent.
-In the SFT stage, the performance and instruction-following abilities of LLMs are largely evoked by carefully constructed instruction datasets Sanh et al.
+(2023a) the techniques used in the construction of their pretraining dataset, leaving the reasons and effects of choosing specific data management strategies absent.
+In the SFT stage, LLMs’ performance and instruction-following abilities are primarily evoked by carefully constructed instruction datasets Sanh et al.
 (2022); Ouyang et al.
 (2022).
-Although a handful of instruction datasets/benchmarks have been proposed with human crowd sourcing Wang et al.
-(2022); Köpf et al.
-(2023), self-instruct Wang et al.
-(2023c); Taori et al.
-(2023) or collection of existing datasets Si et al.
+Although a handful of instruction datasets/benchmarks have been proposed Wang et al.
+(2022, 2023c); Taori et al.
 (2023); Anand et al.
-(2023), practitioners still find it confusing about the effect of instruction datasets on the performance of fine-tuned LLMs, leading to difficulties in choosing proper data management strategies in LLM SFT practices.
-{forest}
-forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=center, font=, rectangle, draw=hidden-draw, rounded corners, align=left, text centered, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=13em,font=,, where level=2text width=14em,font=,, where level=3text width=11em,font=,, [ Data Management, ver [ Pretraining (§2), ver [ Data Quantity (§2.1) [ Scaling Laws [ Kaplan et al.
-(2020), DeepMind Chinchilla Scaling Law Hoffmann et al.
-(2022), leaf, text width=33em ] ] [ Data Repetition [ Villalobos et al.
-(2022), Muennighoff et al.
-(2023), Hernandez et al.
-(2022),
-Xue et al. (2023), D4 Tirumala et al. (2023), leaf, text width=33em ] ] ] [ Data Quality (§2.2) [ Deduplication [ Lee et al. (2021), Kandpal et al. (2022), Silcock et al. (2022),
-SemDeDup Abbas et al. (2023), leaf, text width=33em ] ] [ Quality Filtering [ Gao (2021), Kreutzer et al. (2022), Gunasekar et al. (2023), Li et al. (2023b),
-RefinedWeb Penedo et al. (2023), Marion et al. (2023), Longpre et al. (2023b),
-MiniPileKaddour (2023), leaf, text width=33em ] ] [ Toxicity Filtering [ Luccioni and Viviano (2021), Xu et al. (2021), Welbl et al. (2021),
-Longpre et al. (2023b), leaf, text width=33em ] ] [ Social Bias [ Dodge et al. (2021), Meade et al. (2022), Gururangan et al. (2022)
-Feng et al.
-(2023), leaf, text width=33em ] ] [ Diversity & Age [ Lee et al.
-(2023a), D2 Pruning Maharana et al.
-(2023), Longpre et al.
-(2023b), leaf, text width=33em ] ] ] [ Domain Composition (§2.3) [ Longpre et al.
-(2023b), CodeGen2 Nijkamp et al.
-(2023),
-SlimPajama-DC Shen et al. (2023), DSIR Xie et al. (2023b),
-DoReMi Xie et al.
-(2023a), DoGE Fan et al.
-(2023), leaf, text width=33em ] ] [ Data Management Systems (§2.4) [ Data-Juicer Chen et al.
-(2023a), Oasis Zhou et al.
-(2023c), leaf, text width=33em ] ] ] [ Supervised Fine-Tuning (§3), ver [ Data Quantity (§3.1) [ Ji et al.
-(2023), LIMA Zhou et al.
-(2023a), Yuan et al.
-(2023),
-Chen et al. (2023b), DMT Dong et al. (2023), Song et al. (2023), leaf, text width=33em ] ] [ Data Quality (§3.2) [ Instruction Quality [ INSTRUCTEVAL Chia et al. (2023), LIMA Zhou et al. (2023a),
-Ding et al. (2023), Wang et al. (2023d), Li et al. (2023a),
-Instruction Mining Cao et al. (2023), Self-Refine Madaan et al. (2023),
-SELF Lu et al. (2023a), SelFee Ye et al. (2023), leaf, text width=33em ] ] [ Instruction Diversity [ UltraChat Ding et al. (2023), LIMA Zhou et al. (2023a),
-Alpaca Taori et al. (2023), #InsTag Lu et al. (2023b),
-Self-Instruct Wang et al. (2023c), Explore-Instruct Wan et al. (2023), leaf, text width=33em ] ] [ Instruction Complexity [ #InsTag Lu et al. (2023b), WizardLM Xu et al. (2023a),
-WizardCoder Luo et al. (2023), Orca Mukherjee et al. (2023),
-Tree-Instruct Zhao et al. (2023b), CELLO He et al. (2023),
-FollowBench Jiang et al. (2023), leaf, text width=33em ] ] [ Prompt Design [ Mishra et al. (2022), Khashabi et al. (2022), Gonen et al. (2022),
-Yin et al. (2023b), Kung and Peng (2023), UIT Liang et al. (2023),
-Weber et al.
-(2023), Gudibande et al.
-(2023), Song et al.
-(2023), leaf, text width=33em ] ] ] [ Task Composition (§3.3) [ Wei et al.
-(2021), Wang et al.
-(2022), Sanh et al.
-(2022), Chung et al.
-(2022),
-Flan 2022 Longpre et al. (2023a), ELM Jang et al. (2023), Chen et al. (2023b)
-DMT Dong et al.
-(2023), OPT-IML Iyer et al.
-(2022), Tulu Wang et al.
-(2023b), leaf, text width=34em ] ] [ Data-Efficient Learning (§3.4) [ AlShikh et al.
-(2023), Attendu and Corbeil (2023), Ivison et al.
-(2023),
-Instrucion Mining Cao et al. (2023), AlpaGasus Chen et al. (2023c),
-OpenChat Wang et al. (2023a), DiverseEvol Wu et al. (2023),
-Dynosaur Yin et al. (2023a), MAmmoTH Yue et al. (2023),
-DMT Dong et al. (2023), LoBaSS Zhou et al. (2023b),
-Xu et al. (2023b), Data-Juicer Chen et al. (2023a), leaf, text width=34em ] ] ] ]
-Figure 1: Taxonomy of research in data management for pretraining and supervised fine-tuning of Large Language Models (LLM).
-To address these challenges, it is necessary to conduct a systematic analysis on LLM data management, including the rationale behind management strategy selection and its consequential effect, the evaluation of curated training datasets, and the pursuit of improved strategies.
-Therefore, this survey aims to provide a comprehensive overview of current research in LLM data management, as shown in Figure 1.
-In Section 2, we focus on LLM pretraining data management, including the research on data quantity, data quality, domain composition, and data management systems.
-In Section 3, we discuss the data quantity, data quality, task composition, and data-efficient learning in the SFT stage of LLMs.
-In Section 4, looking into the future, we present the existing challenges and promising future directions in training data management for LLMs.
-Through this survey, we are devoted to offering a guiding resource to practitioners attempting to build powerful LLMs with effective and efficient data management practices.
+(2023), practitioners still find it confusing about the effects of instruction datasets on the performance of fine-tuned LLMs, leading to difficulties in choosing proper data management strategies in LLM SFT practices.
+To address the sparsity problem of existing data, collecting data from multimodal source Zhang et al.
+(2023a); Yang et al.
+(2023b) and model synthesis Maini et al.
+(2024); Li et al.
+(2024a) rise as new trends.
+To address these challenges, researchers try to discover and explore the underlying principles of data management.
+With more and more works been proposed to address different aspects, it is necessary to conduct a systematic discussion considering the whole picture.
+This survey aims to provide a comprehensive overview of current research in LLM data management and a guiding resource to practitioners attempting to build powerful LLMs with efficient data management practices.
+In Section 2 and 3, we respectively discuss current research in the pretraining and SFT stages of LLMs, covering multiple aspects in data management like domain/task composition, data quality, data quantity, etc., as shown in Figure 3.
+However, there still lacks a well-established and acknowledged general data management pipeline.
+Hence, We hope our work can inspire future research to establish and analyze such general pipelines.
+With the vision that the development of data management should keep pace with that of LLMs’ abilitites, we present more existing challenges and promising future directions in Section 4.
+Figure 1: Data management pipelines for the pretraining and supervised fine-tuning of Large Language Models.
 2 Pretraining of LLM
-Data management is found to be important in the pretraining stage of many prominent LLMs OpenAI (2023); Touvron et al.
-(2023a); Wei et al.
+Data management is found to be important in the pretraining stage of many prominent LLMs OpenAI (2023); Touvron et al. (2023a); Wei et al. (2022).
+In this section, we will discuss works trying to explore data management in the pretraining stage of LLMs, including domain composition, data quantity and data quality, as shown in Figure 1(a).
+Strategies adopted by prominent pretrained models are listed in Table 1.
+2.1 Domain Composition
+Publicly available pretraining datasets, like the Pile Gao et al.
+(2020), usually contain mixtures of data collected from multiple sources and domains.
+Many prominent models Du et al.
+(2022); Gao et al.
+(2023); Zhang et al.
+(2023a) are also trained on a mixture of data from different domains.
+Figure 2 summarizes the revealed domain mixture ratios in the pretraining datasets of prominent models.
+Figure 2: The domain composition of prominent Large Language Models.
+Early pretraining corpus mostly contain data with high diversity (Web and Wiki).
+With recent emphasis on the data quality and the requirement for advanced abilities, high quality text (Books and academic text) are integrated.
+Most recently, with improved importance of Coding LLMs and the essential finding that code-based pretraining can enhance reasoning capability of LLMs Liang et al.
+(2022); Guo et al.
+(2024), domain data like code and math take up higher ratio of the total pretraining data.
+A trend can be concluded that more and more domains are included to pretrain LLMs with more various and powerful abilities.
+The benefits of multi-domain composition are also studied in a recent study Longpre et al.
+(2023b).
+Proper domain mixture ratio is also important in the pretraining of LLMs.
+Early attempts usually found the ratio by elaborated experiments and intuitions Gao et al.
+(2020); Du et al.
+(2022); Thoppilan et al.
 (2022).
-Understanding the effects of these data management strategies is also crucial for building strong LLMs.
-However, most of the prominent LLMs do not report their data management procedures or only report the strategies they adopted without discussing the reason for choosing the specific strategies.
-Thus, some researchers try to disclose the working scheme of data management in the pretraining stage of LLMs.
-In this section, we first review the research studying pretraining dataset scaling law with/without data repetition.
-Then, data quality regarding deduplication, quality filtering, toxicity filtering, social bias, and data diversity and age are explored.
-After that, domain composition and domain re-weighting methods are discussed.
-Finally, two recently proposed data management systems are introduced to implement the pretraining data management pipelines.
+Recently, domain generalization techniques are leveraged to automatically assign domain weights to form a suitable target distribution, such as importance resampling Xie et al.
+(2023b) and Group Distributionally Robust Optimization Xie et al.
+(2023a).
+Contribution of each domain measured via gradients is also adopted to reweight domains Fan et al.
+(2023).
+Xia et al.
+(2023) assign batch-level weights dynamically based on varying losses.
+Ye et al.
+(2024) propose data mixing laws to predict model performance with different mixing ratios.
+Although proper domain composition is broadly acknowledged as beneficial in the pretraining of LLMs, as discussed previously, some empirical analyses arrive at different conclusions and leave open questions for future research.
+For example, Longpre et al.
+(2023b) claim that the inclusion of diverse web domains may perform better than specific mixtures in certain tasks.
+Nijkamp et al.
+(2023) study programming and natural language mixtures and find that models trained with mixtures do not perform better than but closely to domain-matched models given the same computing budget.
 2.2 Data Quantity
-The amount of data required for efficient pretraining of LLMs is an ongoing research topic in NLP communities.
-Scaling laws are proposed to depict the relationship between model size and training dataset size.
-According to the proposed scaling laws Kaplan et al.
-(2020); Hoffmann et al.
-(2022), with model size continuously increasing, the demand for more training data will also increase consistently.
-Thus, the exhaustion of text data draws researchers’ attention to data repetition in the LLMs’ pretraining.
+It is recognized that the pretraining of LLMs needs large amounts of data.
+Scaling laws are proposed to depict the relationships between data quantity and model size.
+Repeatedly training on data is also studied due to data exhaustion.
 2.2.1 Scaling Laws
@@
 Kaplan et al.
-(2020) use Transformers and cross-entropy loss to study the empirical scaling laws for language model performance and find the model performance has a power-law relationship with training dataset size or model size, respectively, when not bottlenecked by each other and the training compute budget:
-$L(D)=\big{(}\frac{D_{c}}{D}\big{)}^{\alpha_{D}};\alpha_{D}\sim 0.095,D_{c}\sim 5.4\times 10^{13}$ (1)
-$L(N)=\big{(}\frac{N_{c}}{N}\big{)}^{\alpha_{N}};\alpha_{N}\sim 0.076,N_{c}\sim 8.8\times 10^{13}$ (2)
-where $L$ is the test loss, $D$ is the number of training tokens, $N$ is the number of model parameters, $\alpha_{D}$ and $\alpha_{N}$ is the power-law component for the scaling of $D$ and $N$ respectively, and $D_{c}$ and $N_{c}$ is constant numbers measured in the number of training tokens and non-embedding parameters respectively whose precise numerical value depends on vocabulary size and tokenization and does not have fundamental meaning.
-Combining Equation 1 and  2, they derive Equation 3 to depict the dependence between model size $N$ and dataset size $D$.
-Fitting Equation 3, they get $\alpha_{D}=0.103$, $\alpha_{N}=0.076$, $N_{c}=6.4\times 10^{13}$ and $D_{c}=1.8\times 10^{13}$ and conclude that model performance improves predictably as long as the model size and training dataset size are scaled up simultaneously but encounters overfitting if either of them is fixed while the other increases.
-Given fixed compute budget $C$, they analyze the optimal allocation of $D_{opt}\sim C^{0.27}$ and $N_{opt}\sim C^{0.73}$, showing that the model size should increase faster than the training dataset size.
+(2020) find that the language model loss has a power-law relationship with training dataset size or model size, respectively, when not bottlenecked by each other and the training computing budget.
+They further depict the dependence between model size and training dataset size as:
 $L(N,D)=\big{[}\big{(}\frac{N_{c}}{N}\big{)}^{\frac{\alpha_{N}}{\alpha_{D}}}+\frac{D_{c}}{D}\big{]}^{\alpha_{D}}$ (1)
-Following the power-law relationship proposed by Kaplan et al. (2020), Hoffmann et al. (2022) conduct experiments on much larger language models and arrive at a new scaling law:
+where $L$ is the language model test loss, $D$ is the number of training tokens, $N$ is the number of model parameters, $\alpha_{D}$ and $\alpha_{N}$ are the power-law components for the scaling of $D$ and $N$, respectively, and $D_{c}$ and $N_{c}$ are constant numbers  [^1] The precise numerical values of $D_{c}$ and $N_{c}$ depend on vocabulary size and tokenization and do not have fundamental meaning..
+Fitting Equation 1, they conclude that model loss decreases predictably as long as the model size and training dataset size are scaled up simultaneously.
+Still, overfitting will happen if either of them is fixed while the other increases.
+Given fixed computing budget $C$, they analyze the optimal allocation of $D_{opt}\sim C^{0.27}$ and $N_{opt}\sim C^{0.73}$, showing that the model size should increase faster than the training dataset size.
+Following Kaplan et al. (2020), Hoffmann et al. (2022) conduct experiments on much larger language models and arrive at a new scaling law, usually called as Chinchilla Scaling Law:
 $L(N,D)=E+\frac{A}{N^{\alpha}}+\frac{B}{D^{\beta}}$ (2)
-where $E=1.69$, $A=406.4$, $B=410.7$, $\alpha=0.34$ and $\beta=0.28$.
-They also analyze the optimal allocation of $N_{opt}\sim C^{0.46}$ and $D_{opt}\sim C^{0.54}$.
-Hence, they draw a different conclusion that model size and pretraining dataset size should scale at roughly the same rate with a larger computing budget.
+where they empirically fit $E=1.69$, $A=406.4$, $B=410.7$, $\alpha=0.34$ and $\beta=0.28$.
+The optimal allocation of $D_{opt}$ and $N_{opt}$ are also analyzed as $D_{opt}\sim C^{0.54}$ and $N_{opt}\sim C^{0.46}$.
+Hence, they draw a different conclusion that model and training dataset sizes should scale roughly at the same rate with a larger computing budget.
+Su et al.
+(2024) dig deeper into Kaplan’s scaling laws and provide more detailed instructions to fit the constants.
 2.2.2 Data Repetition
@@
 (2020) and Hoffmann et al.
-(2022) focus on scaling law with unique data trained only for one epoch, Hernandez et al.
-(2022) address the issue about text overlap in the training dataset and study the scaling law with a small fraction of repeated data.
-They observe a strong double descent phenomenon Nakkiran et al.
-(2021) caused by repeated data, where a peak of test loss appears in the middle range of repetition frequency, i.e., the number of epochs trained on repeated data.
-They also show that repeated data can cause a divergence from power-scaling law Kaplan et al.
-(2020) on model sizes larger than 100M parameters.
-According to the scaling law, more training data is required as the model size grows, raising concerns about the exhaustion of high-quality training data Villalobos et al.
+(2022) both focus on scaling laws with unique data trained only for one epoch,  Hernandez et al.
+(2022) study the scaling laws with a small fraction of repeated data in the training dataset and find that the text overlap may be harmful to model performance, causing a divergence from Kaplan’s scaling law on model size larger than 100M parameters.
+With the models grow larger and larger, data has becoming more and more demanding, raising concerns about the exhaustion of high-quality training data Villalobos et al.
 (2022); Hoffmann et al.
 (2022).
-One straightforward way to overcome this issue is to train on data repeatedly.
-However, repeated data subsets may lead to performance degradation as claimed by Hernandez et al.
+Addressing these concerns, several works study the consequence of repeatedly pretraining on the whole datasets for multiple epochs.
+Scaling law on repeated training is proposed to depict the diminishing of returns with more repetition and larger model sizes Muennighoff et al.
+(2023) and shows a multi-epoch degradation phenomenon Xue et al.
 (2023).
-Motivated by this contradiction, several works study the consequence of pretraining on the whole datasets repeatedly for multiple epochs.
+Further analysis digs out that dataset size, model parameters, and training objectives are the key factors to this phenomenon, and classic regularization techniques may not be helpful, except for dropout Xue et al.
+(2023).
+There are still positive results in the research of data repetition.
 Muennighoff et al.
-(2023) find that with constrained data and fixed compute budgets, repeatedly training on the whole dataset up to 4 epochs only causes trivial harm to test loss compared to training on unique new data.
-They also propose a scaling law on repeated training depicting the diminishing of returns with more repetition and larger model sizes.
-Xue et al.
-(2023) also observe a multi-epoch degradation in model performance and find that dataset size, model parameters, and training objectives are the key factors to this phenomenon.
-They further find that commonly used regularization techniques are not helpful in alleviating multi-epoch degradation, except for dropout.
+(2023) find that repeatedly training on the whole dataset up to 4 epochs only causes trivial harm to test loss compared to training on unique new data.
 Instead of simply repeating over the whole dataset, Tirumala et al.
-(2023) show that repeatedly training on carefully selected data can outperform that on randomly selected new data, whilst repeatedly training on randomly selected data cannot, suggesting a feasible way of repeating on intelligently selected data.
+(2023) show that repeatedly training on carefully selected data can outperform that on randomly selected new data, suggesting a feasible way of repeating on intelligently selected data.
+Recently, pretraining with mixed real and synthesized data is adopted to meet the data exhaustion challenge Javaheripi and Bubeck (2023); Meta (2024).
+It is also gaining more an more attention and develops into a new trend as data synthesize.
 2.3 Data Quality
-High-quality data is crucial in the training of machine learning tasks according to previous studies Jain et al.
-(2020); Gupta et al.
-(2021).
-In the pretraining of LLMs, quality assurance techniques are also adopted and usually form a data management pipeline Rae et al.
+In the pretraining of LLMs, Quality control techniques of the pretraining datasets usually form an order Rae et al.
 (2021); Nguyen et al.
 (2023); Tirumala et al.
-(2023), including deduplication, quality filtering, and toxicity filtering.
-Other aspects like social bias, data diversity, and data age are also influential topics in the research community.
+(2023); Gan et al.
+(2023), namely quality filtering, deduplication and toxicity filtering.
+Data diversity and age are also explored.
+2.3.1 Quality Filtering
+Public datasets like Common Crawl  [^2] https://commoncrawl.org/, a large text corpus contains raw web page data, metadata extracts, and text extracts.
+and multilingual datasets Kreutzer et al.
+(2022) usually contain low-quality data that hampers the training of LLMs.
+Hence, existing works usually perform quality filtering using hand-crafted heuristics Yang et al.
+(2019); Raffel et al.
+(2020); Nijkamp et al.
+(2022), a trained classifier Brown et al.
+(2020); Gao et al.
+(2020); Du et al.
+(2022); Touvron et al.
+(2023a); Wettig et al.
+(2024), metric thresholding Wenzek et al.
+(2020); Muennighoff et al.
+(2023) or combinations of these techniques.
+Besides instance-level filtering, embedding clustering is also adopted to filter one cluster at a time Kaddour (2023).
+Despite the reduction of data quantity, quality filtering is usually proven to be beneficial in model performance improvement Longpre et al.
+(2023b).
+Several carefully filtered high-quality datasets are proposed to train lightweight LLMs with outstanding performances Gunasekar et al.
+(2023); Li et al.
+(2023d); Javaheripi and Bubeck (2023); Penedo et al.
+(2023).
+However, Gao (2021) finds that aggressive filtering might lead to performance degradation on a wide range of tasks for GPT-like LLMs due to the poor representativity of the filtering proxy objectives.
+To address this issue, Marion et al.
+(2023) comprehensively examine different data quality estimators and find that pruning datasets based on perplexity performs better than more complicated techniques like memorization.
+Gan et al.
+(2023) develop data-centric scaling laws and show that improving semantic and grammatical quality is more effective.
+However, there still lacks a well-established and theoretically efficient filtering strategy, leaving room for further exploration.
 2.3.2 Deduplication
-Deduplication is widely used in data management procedures of many prominent LLMs and preprocessing of publicly available datasets Brown et al.
-(2020); Workshop et al.
+Deduplication is a necessary step in many LLMs’ pretraining data management procedures and the preprocessing of many publicly available datasets Brown et al.
+(2020); Le Scao et al.
 (2023); Touvron et al.
@@
 Lee et al.
-(2021) use N-gram similarity with MinHash Broder (1997) to detect duplications in training datasets and find that deduplication is beneficial in memorization mitigation, train-test overlap avoidance, and training efficiency improvement while keeping model perplexity.
+(2021) find that deduplication is beneficial in memorization mitigation, train-test overlap avoidance, and training efficiency improvement while keeping model perplexity.
 Kandpal et al.
@@
 (2021).
-Silcock et al.
-(2022) compare it with two neural approaches (a contrastively trained bi-encoder and a "re-rank" style approach combining a bi- and cross-encoder) and conclude that neural approaches can significantly outperform traditional N-gram-and-hashing methods.
-Abbas et al.
-(2023) propose SemDeDup to remove semantic duplicates that lie closely in the pre-trained model embedding space and apply clustering to reduce the searching computation.
-2.2.2 Quality Filtering
-Quality filtering is another key step in constructing a high-quality pretraining dataset, because public datasets like Common Crawl  [^1] https://commoncrawl.org/, a large text corpus contains raw web page data, metadata extracts, and text extracts.
-and multilingual datasets Kreutzer et al.
-(2022) usually contain low-quality data that hampers the training of LLMs.
-Existing works usually perform quality filtering using a classifier Brown et al.
-(2020); Gao et al.
-(2020); Du et al.
-(2022); Touvron et al.
-(2023a), hand-crafted heuristics Yang et al.
-(2019); Raffel et al.
-(2020); Nijkamp et al.
-(2022) or threshold filtering using criterion like perplexity Wenzek et al.
-(2020); Muennighoff et al.
+It can operate at line-level Touvron et al.
+(2023a), document-level Hoffmann et al.
+(2022); Li et al.
+(2022b) or combinations of them.
+Recently, neural models are experimentally proven to outperform traditional N-gram-and-hashing methods Silcock et al.
 (2022).
-Kaddour (2023) construct a subset of the Pile Gao et al.
-(2020) called MiniPile by filtering out low-quality embedding clusters.
-Quality filtering is usually proven to be beneficial in model performance improvement Longpre et al.
-(2023b), despite the reduction of training data quantity and variety.
-Light-weight language model phi-1 Gunasekar et al.
-(2023) and phi-1.5 Li et al.
-(2023b) with 1.3B parameters trained on carefully selected high-quality data and synthetically generated data show outstanding performance on coding and commonsense reasoning tasks.
-Recently, Microsoft published a new 2.7B language model phi-2 Javaheripi and Bubeck (2023) with better language understanding and logic reasoning abilities.
-Penedo et al.
-(2023) construct the RefinedWeb dataset consisting of properly filtered and deduplicated high-quality web data, outperforming models trained on the Pile Gao et al.
-(2020).
-However, Gao (2021) finds that aggressive filtering can lead to performance degradation on a wide range of tasks for GPT-like LLMs because the filtering proxy objectives are not representative enough of the true objective.
-To address this issue, Marion et al.
-(2023) comprehensively examines the combinations of three data quality estimators (perplexity, Error L2-Norm (EL2N), and memorization factor), three proportions of remaining data (top, middle, and bottom proportion ordered by corresponding estimators) and different percentage of remaining data through data pruning.
-Surprisingly, they find that pruning datasets based on perplexity and retaining the middle proportion of data performs far better than more complicated techniques like memorization.
-However, no combinations of pruning strategies achieve consistently high performance.
+Addressing semantic deduplication, Abbas et al.
+(2023) propose SemDeDup to remove semantic duplicates that lie closely in the pretrained model’s embedding space and apply clustering to reduce the searching computation.
 2.3.3 Toxicity Filtering
@@
 (2021) and Welbl et al.
-(2021) both find that training dataset detoxification leads to the marginalization of minority groups like dialects and minority identity mentions.
-2.2.4 Social Bias
-Besides the marginalization of minority groups caused by data detoxifying, several works Kurita et al.
+(2021) find that training dataset detoxification leads to the marginalization of minority groups like dialects and minority identity mentions, posing challenges in building unbiased LLMs.
+2.3.4 Data Diversity
+Some works focus on other aspects of data management in the pretraining stage of LLMs.
+Lee et al.
+(2023a) show that the format diversities of publicly available pretraining datasets are high when measured by Task2Vec diversity coefficient Miranda et al.
+(2022).
+Maharana et al.
+(2023) propose D2 Pruning to balance data diversity and difficulty in data selection by representing datasets as undirected graphs and adopting forward-and-reverse message passing strategy to select a subgraph enveloping both diverse and difficult data samples.
+2.3.5 Data Age
+In current practices, more recent LLMs are usually pretrained using newer data  [^3] https://platform.openai.com/docs/models.
+Some knowledge learned by pretrained LLMs could also be time-sensitive.
+Longpre et al.
+(2023b) study the impact of data age and find that the temporal shift between evaluation and pretraining data will lead to inaccurate performance estimation.
+This temporal misalignment might not be overcome by fine-tuning, especially for larger models.
+2.4 Relations Among Domain Composition, Data Quantity and Data Quality
+Recently, several scaling laws are proposed to explore the synergistic effect of different aspects on the pretrained model performance, such as the bivariate model performance prediction regarding data quantity and domain composition ratio Ge et al.
+(2024a), the quality-quantity tradeoff under different computing budget Goyal et al.
+(2024), and the positive correlation between data quality and model scale under the same data quantity Bi et al.
+(2024).
+What’s more, Shen et al.
+(2023) emphasize global deduplication to remove overlaps among different domains.
+Longpre et al.
+(2023b) claim that domains with high quality and diversity are more beneficial than other domains.
+3 Supervised Fine-Tuning of LLM
+Based on the general knowledge and capabilities learned in the pretraining stage, supervised fine-tuning (SFT) is proposed to further improve LLMs with instruction-following ability and alignment with human expectations Wei et al.
+(2021); Sanh et al.
+(2022); Ouyang et al.
+(2022).
+Although LLMs fined-tuned with existing instruction datasets have achieved remarkable performance in various NLP tasks, the impacts of instruction data management on fine-tuned models are still under debate.
+The data management process in the SFT stage can be summarized as illustrated in Figure 1(b), including task composition, data quality control, data quantity control and dynamic data-efficient learning.
+Table 2 summarizes the data management practices of prominent fine-tuned LLMs.
+3.1 Task Composition
+Since LLMs have shown surprisingly emergent abilities in handling various NLP tasks, multitask fine-tuning appears to be promising to improve LLMs’ generalization performance on unseen tasks.
+The benefits of increasing the number of tasks in SFT have been experimentally proven on models with different sizes ranging from 3B to 540B parameters Wang et al.
+(2022); Sanh et al.
+(2022); Wei et al.
+(2021); Chung et al.
+(2022).
+With the scaling of tasks, the mixture ratio of data targeting different tasks is also found to be critical and usually decided by experiments and intuitions Iyer et al.
+(2022); Longpre et al.
+(2023a).
+To enable LLMs to solve targeted tasks with specific skills, representation similarity Ivison et al.
+(2023); Lee et al.
+(2024) and gradient similarity Xia et al.
+(2024) is proposed to select relevant multitask subsets.
+However, conflicts might exist among the many tasks.
+Dong et al.
+(2023) focus on task composition among mathematical reasoning, code generation, and general human-aligning abilities.
+They find that model abilities are improved when the mixed data amount is small but decreased otherwise.
+The negative impact of large amount mixing data might lie in the similarity degree of data format and data distribution among different SFT tasks.
+Wang et al.
+(2023b) also experimentally show that different instruction datasets may correspond to different specific abilities.
+And winning across all evaluations using a single dataset or combination seems to be challenging.
+Divergent from compositing multiple tasks, some works claim that integration of LLMs tuned on single task data can outperform one LLM tuned on multiple tasks Jang et al.
+(2023); Chen et al.
+(2023b).
+But fine-tuning more task-specific LLMs also means more resource consumption.
+How to efficiently equip LLMs with the ability to solve multiple tasks still demands more exploration.
+3.2 Data Quality
+Data quality is always a focal point in the SFT of LLMs, addressing instruction quality, diversity, and complexity.
+Here, we focus more on managing and analyzing existing instruction data instead of instruction generation methods discussed in previous surveys Zhang et al.
+(2023b); Wang et al.
+(2023e).
+3.2.1 Instruction Quality
+Many researchers have found that the quality of instruction data is one of the most important factors in improving model performance Chia et al.
+(2023); Zhou et al.
+(2023a); Ding et al.
+(2023).
+During the construction of instruction dataset, there is usually a filtering step to select high-quality instructions generated by models.
+Heuristic- and model-based natural language indicators like perplexity and uncertainty are commonly adopted filtering criteria Wang et al.
+(2023d); Cao et al.
+(2023); Bhatt et al.
+(2024).
+What’s more, losses Zhou et al.
+(2023b); Li et al.
+(2023b, 2024b) and output probabilities Li et al.
+(2023a, e); Chen and Mueller (2024); He et al.
+(2024b); Liu et al.
+(2024) of LLMs are adopted to compute more complex scores for data selection.
+Popular searching approaches like BlendSearch Wang et al.
+(2020) are also leveraged to find high-quality instructions satisfying the criteria Cao et al.
+(2023).
+In addition, LLMs are also queried to directly evaluate the quality of instructions.
+Fine-tuned LLMs are prompted to assign quality scores Li et al.
+(2023c) or provide self-feedback Lu et al.
+(2023a); Madaan et al.
+(2023) to their own responses to iteratively improve model prediction.
+Strong LLMs like ChatGPT Ye et al.
+(2023); Chen et al.
+(2023c); Li et al.
+(2023a) or reward models Du et al.
+(2023) are also adopted as quality judges during instruction data filtering.
+Recently, weak-to-strong strategy is introduced to select high-quality data with smaller and weaker models Li et al.
+(2024c); Yang et al.
+(2024); Mekala et al.
+(2024).
+3.2.2 Instruction Diversity
+The intention and semantic diversity of instructions is another important factor that has shown positive effects on model performance improvement and robustness Zhou et al.
+(2023a); Ding et al.
+(2023); Taori et al.
+(2023); Bukharin and Zhao (2023).
+However, there is no well-acknowledged measurement to quantitatively indicate the diversity of an instruction dataset.
+#InsTag Lu et al.
+(2023b) propose to measure instruction diversity using fine-grained tags generated by ChatGPT  [^4] https://chatgpt.openai.com/.
+Specifically, it quantifies instruction diversity as the unique tag coverage rate in the overall tag set.
+To maintain both diversity and data-efficiency in the instruction datasets, Rouge-L similarity Wang et al.
+(2023c), embedding distance Wu et al.
+(2023); Bukharin and Zhao (2023); Huang et al.
+(2024) and scoring models Ge et al.
+(2024b) are proposed to select instructions that are different from each other in literal, semantic and human-aligning level.
+Due to data constraints, diversity can be challenging in some domain-specific tasks.
+Thus, Wan et al.
+(2023) propose to enlarge the data coverage through active searching variations and possibilities of instructions using LLMs.
+3.2.3 Instruction Complexity
+Instruction complexity is found to be crucial in developing LLMs with complex instruction-following and reasoning abilities Xu et al.
+(2023a); Luo et al.
+(2023b); Mukherjee et al.
+(2023); He et al.
+(2024a).
+Several works endeavor to quantify and evaluate instruction complexity.
+Using aforementioned tags, #InsTag Lu et al.
+(2023b) quantifies complexity as the average tag number assigned to each query in a dataset.
+He et al.
+(2023) evaluate complex instruction with eight features addressing the length, contents, and formats of input texts and task descriptions.
+It is also empirically showed that complexity enhancement is necessary for performance improvement Zhao et al.
+(2023b).
+To increase the instruction complexity in SFT datasets, some works propose to incrementally augment existing instructions by adding nodes to semantic tree Zhao et al.
+(2023b) or performing operations such as increasing reasoning, adding constraints, in-breadth evolving, deepening, and so on Xu et al.
+(2023a); Luo et al.
+(2023b); Jiang et al.
+(2023b); Sun et al.
+(2024a).
+3.3 Data Quantity
+Different with the acknowledged scaling laws of pretraining data, explorations of the relationship between scaling instruction data quantity and fine-tuned model performance diverge in two directions.
+In the earlier stage, researchers follow the observations in the pretraining of LLMs and argue that scaling up the instruction data quantity is crucial for success Wei et al.
+(2021); Sanh et al.
+(2022).
+Recently, more works claim that data quality is more important than data quantity in the SFT of LLMs, and propose to scaling down the instruction datasets with limited high-quality data Zhou et al.
+(2023a); Chen et al.
+(2023b).
+However, Zhang et al.
+(2024) propose a power-based multiplicative joint scaling law, showing that increased fine-tuning data could lead to improved model performance after achieving good results with limited data.
+Addressing this conflict, several works attempt to analyze the scaling patterns for different tasks or different model abilities.
+A consensus of these works is that different abilities have different scaling patterns and develop at different paces.
+Dong et al.
+(2023) find that general ability can be enhanced with about 1,000 samples and improves slowly after then, while mathematical reasoning and code generation improve consistently with the increasing of instruction data amount.
+Similarly, Yuan et al.
+(2023) observe a log-linear relation between instruction data amount and models’ mathematical reasoning performance, but stronger pretrained models improve less with more instruction data.
+Surprisingly, the empirical study of Ji et al.
+(2023) on 12 major real-world online user cases draws to an exactly opposite point.
+Song et al.
+(2023) also show that some abilities have completely different patterns from others.
+3.4 Dynamic Data-Efficient Learning
+While works discussed above focus more on the static management of instruction datasets without interaction with model fine-tuning, some works try to combine data selection with model fine-tuning, achieving data-efficient learning in a dynamic way.
+Training affects data.
+Some works propose to dynamically change the datasets along with the fine-tuning process.
+Attendu and Corbeil (2023) propose a dynamic data pruning method that periodically filters out unimportant examples during SFT using extended versions of EL2N metric Paul et al.
+(2021); Fayyaz et al.
+(2022).
+AlShikh et al.
+(2023) predict the responses as "answer-like or not" by a binary classifier, in order to measure LLMs’ instruction-following ability and serve as an early-stopping criterion.
+Kung et al.
+(2023) conduct active task searching to select informative tasks based on prompt uncertainty and fine-tune in a loop.
+Data affects training.
+Instead of manipulating instruction datasets, some works propose special training strategies to accommodate the datasets.
+To mitigate forgetting and negative task impact, Yin et al.
+(2023a) and Wang et al.
+(2024) treat task selection as a replay strategy in continual learning scenarios; DMT Dong et al.
+(2023) learns specialized and general abilities sequentially while keeping a small proportion of specialized data.
+To efficiently learn mixed-quality data acquired from LLMs with different level of abilities, OpenChat Wang et al.
+(2023a) proposes C-RLFT strategy that considers different data sources as coarse-grained reward labels; Xu et al.
+(2023b), Sun et al.
+(2024a) and Kim and Lee (2024) propose to make the model progressively learn from easy to hard, respectively regarding different data quality, instruction complexity and task hardness.
+3.5 Relations Among Task composition, Data Quality and Data Quantity
+Similar as in the pretraining stage, different aspects of supervised fine-tuning data management can affect model performance jointly.
+Lu et al.
+(2023b) analyze popular open-set SFT datasets using #InsTag and show that larger dataset sizes tend to be more diverse and induce higher performance.
+Current research on data selection tends to uniformly consider instruction quality and diversity Bukharin and Zhao (2023); Xu et al.
+(2023c).
+Since different model abilities have different scaling patterns as discussed in Section 3.3, more efficient task composition strategies are required to build stronger multi-task LLMs.
+In summary, we provide a list of takeaways in Appendix A. Some other aspects of data management are discussed in Appendix B.
+4 Challenges and Future Directions
+The exploration of data management and its impact on LLM pretraining and SFT is still an ongoing task.
+In this section, we point out several challenges and corresponding future directions in the research of training data management for LLMs.
+General data management framework
+Although data management systems are proposed to compose various data recipes in either the pretraining or SFT stage of LLM Chen et al.
+(2023a); Zhou et al.
+(2023c); Sun et al.
+(2024b), practitioners still need to spend efforts on organizing suitable datasets.
+A well-established general data management framework suitable for a broad range of applications is an urgent and worthy future direction in developing and promoting LLMs.
+Beyond that, a more autonomous data management system is also needed to greatly save human efforts.
+To build such systems, LLMs might be leveraged and serve as different roles such as quality examinator, data augmentor, and so on.
+Data debiasing and detoxifying
+Current pretraining corpora and instruction datasets might contain harmful information and social biases, which lead to negative social impacts and undesirable model behavior.
+With the application of LLMs keeps extending to more demanding fields, the fairness and harmlessness of LLMs will become more and more innegligible.
+Hence, as one way to build ideal LLMs without biases and harmful output, debiasing and detoxifying of pretraining and instruction data is an important research direction.
+Multimodal data management
+Current research in data management mostly focuses on natural language processing.
+With the application of LLMs extending to modalities like vision, audio, etc., it is necessary to see the impacts of multimodal data management on the performance of fine-tuned multimodal LLMs.
+Data management for LLM self-exploration
+The ability to actively explore the unknown environment and tasks is one of the future perspectives in LLM development.
+Learning from large-scale interaction data requires efficient data management system to construct suitable datasets.
+Efficient filtering for synthesized data
+As data annotation requires intensive human labors and existing data will be exhausted, automatically synthesizing new data using LLMs is newly proposed as a promising solution Maini et al.
+(2024); Li et al.
+(2024a).
+In this process, efficient filtering for synthesized data is required to ensure its quality.
+Fine-grained data ordering
+Some works start to pay attention to the ordering of data in both the pretraining Gan et al.
+(2023); Guo et al.
+(2024) and SFT stage Xu et al.
+(2023b); Yin et al.
+(2023a).
+It is shown that more fine-grained data ordering could be beneficial to model performance improvement.
+Conflicted data separation
+In multi-task fine-tuning, negative impact of mixing data is observed and attribute to conflicts among different task data Dong et al.
+(2023).
+Thus, separating and effectively learning from conflicted data samples is a challenging problem in multi-task learning.
+5 Conclusions
+This paper overviews the training data management of LLMs.
+We discuss the pretraining and supervised fine-tuning stages of LLM successively and summarize the up-to-date research efforts according to the data management process of each stage.
+Finally, we highlight several challenges and future directions for LLM training data management.
+We hope this survey can provide insightful guidance for practitioners and inspire further research in efficient training data management for the development of LLMs.
+Appendix A Takeaways
+In the pretraining stage of LLMs:
+  • The coverage of more domains and proper domain mixture ratio are important.
+Recently, researchers try to automatically find the proper domain mixture weights, which still show room for improvement.
+  • Large amount of data is widely considered critical, and proper data repetition may also bring positive impacts to model performance.
+  • Data quality control is necessary usually form an order, namely quality filtering, deduplication and toxicity filtering..
+However, over-aggressive quality and toxicity filtering may lead to performance degradation and social biases, which is still under-explored.
+  • Data diversity and temporal misalignment also have impacts on model performance, which call for future study.
+In the supervised fine-tuning stage of LLMs:
+  • Multitask fine-tuning is widely adopted nowadays.
+However, conflictions may exist among tasks and hinders the model abilities.
+Hence, dealing with negative task confliction is also calling for better answers.
+Ensembling multiple single-task experts instead of training one multitask model also arises as an new trend.
+  • Quality control are usually achieved through heuristics, human evaluation or LLMs as quality judges.
+Instruction diversity and complexity are also beneficial and enhanced by several works.
+The exploration of more diverse and complex instructions is still an open question.
+  • Works have shown that the SFT of LLM rely more on data quality than data quantity.
+However, digging deeper into the influence of data quantity, some researchers find that the learning of different tasks may require different amount of data.
+  • Instead of keep instruction datasets unchanged during fine-tuning, works propose to adjust the datasets dynamically through fine-tuning.
+Special fine-tuning strategies are also continually shown up to utilize the instruction data more efficiently.
+Appendix B Other Aspects of Data Management For LLMs
+B.1 Social Bias
+Besides the marginalization of minority groups caused by data detoxifying mentioned in Section 2.3.3, several works Kurita et al.
 (2019); Nangia et al.
@@
 partisan leanings and how it is propagated to language models even further to downstream tasks.
-2.2.5 Diversity & Age
-There are also works focusing on other aspects of data management in the pretraining stage of LLMs.
-For example, Lee et al.
-(2023a) show that the format diversities of publicly available pretraining datasets are high when measured by the recently proposed Task2Vec diversity coefficient Miranda et al.
-(2022).
-They also demonstrate the effectiveness of the diversity coefficient by showing its alignment with the number of latent concepts in the text and suggest using it to build more diverse datasets.
-Maharana et al.
-(2023) proposes a novel pruning method D2 Pruning to balance data diversity and difficulty in data selection by representing a dataset as an undirected graph with samples as nodes and difficulty scores as node properties.
-Each sample is connected with its k nearest neighbors weighted by distances in the embedding space.
-Then, a forward and reverse message passing strategy is adopted to select a subgraph enveloping both diverse and difficult data samples.
-Longpre et al.
-(2023b) explore the age of the evaluation dataset and draw conclusions that the temporal shift between evaluation and pretraining data will lead to inaccurate performance estimation and the temporal misalignment might not be overcome by fine-tuning, especially for larger models.
-2.3 Domain Composition
-Public available pretraining datasets usually contain mixtures of data collected from multiple sources and domains.
-For example, the Pile Gao et al.
-(2020) contains web documents from Common Crawl, Wikipedia, Books, and collections from medical, academic, coding and math, legal, and social resources.
-Many prominent models are also trained on a mixture of data from different domains.
-For example, LaMDA Thoppilan et al.
-(2022) is trained on dialogs data from public forums, C4 data, Wikipedia (English), web documents, and code documents from programming-related Q&A sites, tutorials, etc.
-Efforts are made to explore the impact of domain mixtures on the pre-trained model performance.
-Longpre et al.
-(2023b) group the Pile Gao et al.
-(2020) data into nine domains and conduct ablate-one-at-a-time experiments to show the impact of different domains.
-They draw conclusions that the domains with high quality (Books) and high diversity (Web) are broadly helpful, and it is beneficial to include as many data sources as possible even though they are less relevant to the downstream tasks.
-SlimPajama-DC Shen et al.
-(2023) arrives at the same point that including all domains typically achieves better results than deliberately selected domain combinations, given that global deduplication is conducted to remove overlaps among different domain datasets.
-Both Longpre et al.
-(2023b) and Shen et al.
-(2023) agree that specific mixtures may excel in evaluation benchmarks for targeted tasks, but the former claim that the inclusion of diverse web domains may perform better than specific mixtures in certain tasks.
-CodeGen2 Nijkamp et al.
-(2023) studies the impact of mixtures of programming and natural languages on model performance and finds that models trained with mixtures do not perform better than but closely to domain-matched models given the same compute budget.
-Several methods are also proposed to find the proper domain composition weights.
-DSIR Xie et al.
-(2023b) formulates the problem as a distribution matching problem given a large raw unlabeled dataset and some unlabeled target samples.
-Specifically, it leverages the classic importance resampling approach Rubin (1988) and estimates the importance weights using n-gram features and KL reduction.
-Without knowledge of downstream tasks or target distributions, DoReMi Xie et al.
-(2023a) trains a small proxy model using Group Domain Robust Optimization (Group DRO) Oren et al.
-(2019); Sagawa* et al.
-(2020) to generate domain weights.
-It improves model performance on all domains by up-weighting domains with the largest loss gap between the evaluated model and a pre-trained reference model.
-Improved from DoReMi Xie et al.
-(2023a), Fan et al.
-(2023) propose DoGE which reweights training domains to minimize the average validation loss across all training domains or on a specific unseen domain.
-The key component of their generalization objective is a gradient-based generalization estimation function measuring the contribution of each domain to other domains.
-Then, domains contributing higher to learning other domains will receive larger weights, as well as those difficult to learn.
-2.4 Data Management Systems
-Addressing the difficulty in pretraining data management, integrated data management systems are beneficial for LLM practitioners with different demands.
-Chen et al.
-(2023a) provides a data processing system Data-Juicer featuring the generation of diverse data recipes.
-They provide over 50 versatile data management operators and dedicated tools targeting zero-code data processing, low-code customization, and off-the-shelf data processing components.
-A timely feedback loop at multiple development stages of data recipes and LLMs is also supported.
-Zhou et al.
-(2023c) also propose a pretraining data curation and assessment system Oasis, which contains an interactive modular rule filter module, a debiased neural quality filter module, an adaptive document deduplication module, and a holistic data assessment module.
-3 Supervised Fine-Tuning of LLM
-Based on the general knowledge and capabilities learned in the pretraining stage, supervised fine-tuning (SFT) is proposed to further improve LLMs with instruction-following ability and alignment with human expectations Wei et al.
-(2021); Sanh et al.
-(2022); Ouyang et al.
-(2022).
-Many efforts have been made to construct instruction data using human crowd-sourcing Wang et al.
-(2022); Köpf et al.
-(2023), self-instruct Wang et al.
-(2023c); Taori et al.
-(2023) or collection of existing datasets Si et al.
-(2023); Anand et al.
-(2023).
-Although LLMs fined-tuned with existing instruction datasets have achieved remarkable performance in various NLP tasks, the impacts of instruction data management on fine-tuned model performance are still under debate.
-Consistent with previous discussion regarding LLM pretraining, in this section, we summarize the research explorations in LLM SFT into data quantity, data quality (instruction quality, diversity, complexity, prompt design), and task composition.
-Data-efficient SFT is also included to discuss current efforts on efficient SFT from the data aspect.
-3.1 Data Quantity
-The explorations of the relationship between scaling instruction data quantity and fine-tuned model performance diverge in two directions.
-One branch of research focuses on scaling down the instruction data quantity to improve training efficiency.
-For example, LIMA Zhou et al.
-(2023a) carefully curated 1,000 high-quality samples and experimentally justified their hypothesis that only limited instruction tuning data is needed to expose the knowledge and capabilities that the LLM has already acquired during pretraining.
-Chen et al.
-(2023b) observe that maybe a single instruction is sufficient for single task-specific LLM fine-tuning, and 16K samples with 1.9M tokens may be sufficient to train a model specialized in the natural language inference (NLI) task.
-Another branch of research argues that scaling up the instruction data quantity is crucial for success Wei et al.
-(2021); Sanh et al.
-(2022).
-Addressing this conflict, several works attempt to analyze the scaling pattern for different tasks or model abilities.
-Ji et al.
-(2023) conduct an empirical study on 12 major real-world online user cases and show that scaling up the instruction data leads to continuous improvement in tasks such as extraction, classification, closed QA, and summarization while leading to little improvement in tasks such as math, code, and chain-of-thought.
-Disagree with Ji et al.
-(2023), Dong et al.
-(2023) find that general ability can be enhanced with about 1,000 samples and improves slowly after then, while mathematical reasoning and code generation improve consistently with the increasing of data amount.
-Similarly, Yuan et al.
-(2023) observe a log-linear relation between instruction data amount and models’ mathematical reasoning performance, but stronger pre-trained models improve less with larger fine-tuning datasets.
-Song et al.
-(2023) conduct experiments covering ten distinct in-domain abilities and three out-of-domain abilities and show that most abilities are consistent with data scaling.
-Still, each ability develops at different paces during instruction tuning, while some abilities show completely different patterns.
-3.2 Data Quality
-Data quality is always a focal point in the SFT of LLMs, addressing instruction quality, diversity, complexity, and prompt design.
-Here, we focus more on the management and analysis of existing instruction data instead of instruction generation methods discussed in previous surveys Zhang et al.
-(2023b); Wang et al.
-(2023e).
-3.2.1 Instruction Quality
-Many researchers have found that the quality of instruction data is one of the most important factors in improving model performance Chia et al.
-(2023); Zhou et al.
-(2023a); Ding et al.
-(2023).
-During the construction of instruction data, there is usually a filtering step to select high-quality instructions generated by models.
-Wang et al.
-(2023d) use perplexity as the criterion to select the most appropriate instructions from the pool of candidate instructions generated by open-source models.
-Cao et al.
-(2023) propose an automatic data selector InstructionMining to evaluate instruction data quality without human experts’ interventions.
-They first hypothesize that the inference loss of a fine-tuned model on an evaluation set can serve as a proxy for data quality.
-Then, they use a set of natural language indicators to predict the inference loss to save the efforts in actually fine-tuning an LLM, including the number of tokens in tokenized inputs and outputs, reward score Köpf et al.
-(2023), perplexity, Measure of Textual Lexical Diversity McCarthy and Jarvis (2010), distance to approximate i-th nearest neighbors Dong et al.
-(2011) in SentenceBERT Reimers and Gurevych (2019) embedding space, and the naturalness, coherence, and understandability scores provided by UniEval dialogue model Zhong et al.
-(2022).
-Instead of using indicators to filter low-quality instructions, several works Li et al.
-(2023a); Lu et al.
-(2023a); Ye et al.
-(2023); Madaan et al.
-(2023) leverage the power of fine-tuned LLM itself to evaluate the quality of instructions.
-Li et al.
-(2023a) assign quality scores to augmented instructions using the language model and iteratively improve model prediction.
-Similarly, SELF Lu et al.
-(2023a) and Self-Refine Madaan et al.
-(2023) prompts LLM to provide self-feedback on their own responses.
-Strong LLMs like ChatGPT are also adopted as quality judges during the instruction collection process Ye et al.
-(2023).
-3.2.2 Instruction Diversity
-The intention and semantic diversity of instructions is another important factor that has shown a positive effect on model performance improvement Zhou et al.
-(2023a); Ding et al.
-(2023); Taori et al.
-(2023).
-Self-Instruct Wang et al.
-(2023c) adopts ROUGE-L similarity to filter out the newly generated instructions that are too similar to the existing ones.
-To better evaluate the instruction diversity of SFT datasets, #InsTag Lu et al.
-(2023b) is proposed as an open-set fine-grained tagger using ChatGPT  [^2] https://chatgpt.openai.com/.
-Specifically, it first prompts ChatGPT to provide tags for given queries in an open setting, then performs a normalization procedure to deal with the noise in the raw tagging, including low-frequency filtering, rule aggregation, semantic aggregation with embedding clustering, and association aggregation that merges associated tags together.
-With the generated tags, they quantify instruction diversity as the unique tag coverage rate for the overall tag set.
-They also analyze popular open-set SFT datasets and show that larger dataset size tends to be more diverse and induces higher performance.
-Though important, diversity can be challenging in domain-specific tasks due to data constraints.
-Wan et al.
-(2023) propose an approach called Explore-Instruct to enlarge the data coverage through active exploration via LLMs.
-Explore-Instruct starts from representative domain user cases and searches the variations and possibilities by looking ahead into potential fine-grained sub-tasks and backtracking alternative branches in the search space.
-3.2.3 Instruction Complexity
-The complexity of instructions also attracts researchers’ attention, especially in developing LLMs with complex instruction-following and reasoning abilities Xu et al.
-(2023a); Luo et al.
-(2023); Mukherjee et al.
-(2023).
-Several works endeavor to quantify and evaluate instruction complexity.
-Using aforementioned tags, #InsTag Lu et al.
-(2023b) quantifies complexity as the average tag number assigned to queries in a dataset.
-He et al.
-(2023) evaluate complex instruction with eight features, i.e., multi-turn, length, noise, and heterogeneous information for input text, multi-tasking, semantics, formats, and quantity constraints for task description.
-To delve into the exploration of instruction complexity, Zhao et al.
-(2023b) propose Tree-Instruct to enhance the complexity of instruction data controllably.
-It treats the instruction as a semantic tree and constructs new complex instructions by adding nodes to the tree.
-Thus, the complexity of instruction can be controlled by adjusting the number of added nodes.
-Through experiments, they find that increased complexity can lead to continuing performance improvement.
-What’s more, the improvement does not come from the increased number of tokens, as a few complex instructions still outperform diverse but simple instructions under the same token budget.
-Curriculum instruction tuning ranging from easy to difficult might not be as helpful as expected, indicating the necessity of enhancing complexity.
-Another method of increasing complexity is also proposed.
-Evol-Instruct Xu et al.
-(2023a); Luo et al.
-(2023) rewrite instructions step by step with operations such as increasing reasoning, adding constraints, in-breadth evolving, deepening, and complicating input with code and table.
-Similarly, Jiang et al.
-(2023) incrementally augment instructions with constraints on content, situation, style, format and example and propose FollowBench to evaluate LLMs’ constraint following ability.
-3.2.4 Prompt Design
+As addressed in previous research, there is still a large gap between current prominent LLMs and ideal LLMs without social biases.
+Many questions are worth exploring, such as how to mitigate the potential biases in pretraining datasets, the existence of bias in the SFT datasets, and whether it is feasible to reduce social bias through SFT.
+B.2 Prompt Design
 Current instructions are either heuristically designed by human Wang et al.
@@
 (2023).
-However, the same intention and semantic meaning can be phrased into various prompts, and the choice of prompts can cause significant model performance variation Gonen et al.
+The choice of prompts might cause significant model performance variation Gonen et al.
 (2022); Weber et al.
 (2023).
-Thus, what kinds of instruction prompts are better for LLM training might be vital yet have not been fully explored.
 Early attempts include manual reformulation of prompts into the ones easier to follow for language models Mishra et al.
-(2022) and choosing prompts with the lowest perplexity to get the most significant gains in model performance Gonen et al.
+(2022), and choosing prompts with the lowest perplexity to get the most significant gains in model performance Gonen et al.
 (2022).
+Recently, Liang et al.
+(2023) develop a format transfer framework UIT to transfer instructions from different datasets into unified formats automatically.
+Some works focus on studying the impact of prompt phrasing.
 Khashabi et al.
 (2022) surprisingly find that the discretized interpretation of continuous prompts is not always consistent with the discrete prompts describing the same task as heuristically expected.
-To know which parts of the instructions are most important in LLM fine-tuning, Yin et al.
-(2023b) and Kung and Peng (2023) both conduct ablation analysis removing contents in task definitions.
 Yin et al.
@@
 They achieve comparable model performance using the modified task definitions and delusive examples containing incorrect input-output mappings.
-Based on this finding, they cast doubts on the performance gain of fine-tuned models and state that the model may only learn superficial patterns during instruction tuning.
-Addressing the issue of instruction format inconsistency, Liang et al.
-(2023) develop a format transfer framework UIT to transfer instructions from different datasets into unified formats automatically.
+Based on their experiment results, they cast doubts on the performance gain of fine-tuned models and state that the model may only learn superficial patterns during instruction tuning.
 Besides the choice of phrasing, the generation source of prompts is another factor in prompt design.
 Gudibande et al.
-(2023) raise questions on fine-tuning a weaker language model on outputs of a stronger model and find that the imitation model might adapt to mimic the stronger model’s style but not its functionality, indicating the potential challenge of closing the performance gap between open- and close-sourced through imitation instruction tuning.
+(2023) raise questions on fine-tuning a weaker language model on outputs of a stronger model and find that the imitation model might adapt to mimic the stronger model’s style but not its functionality.
 Similarly, Song et al.
 (2023) also observe that human-designed data can outperform synthetically generated data from GPT-4 OpenAI (2023) to a relatively large extent.
-3.3 Task Composition
-Since LLMs have shown surprisingly emergent abilities in handling various NLP tasks, multitask fine-tuning appears to be a promising approach to improve LLMs’ generalization performance on unseen tasks.
-The benefits of increasing the number of tasks in SFT have been experimentally proven on models with different sizes ranging from 3B to 540B parameters Wang et al.
-(2022); Sanh et al.
-(2022); Wei et al.
-(2021); Chung et al.
-(2022).
-Besides the scaling of the number of tasks, the mixture ratio of different instruction benchmarks and task balancing is also found to be critical for effective instruction fine-tuning Iyer et al.
-(2022); Longpre et al.
-(2023a).
-Dong et al.
-(2023) focus on task composition among mathematical reasoning, code generation, and general human-aligning abilities.
-Compared with individual source data, They find that model abilities are improved when the mixed data amount is small but decreased when the mixed data amount is large.
-The results indicate that conflicts exist among abilities with larger amounts of mixed data.
-To further explain the conflicts, they vary the ratio of general and specialized data and conclude that the impact of data ratio might lie in the similarity degree of data format and data distribution among different SFT tasks.
-Different from compositing multiple tasks together, some works claim that LLM tuned on a single task data can outperform LLM tuned on multiple tasks Jang et al.
-(2023); Chen et al.
-(2023b).
-Jang et al.
-(2023) state that training expert LLMs to form an expert library can outperform multitask language models by avoiding negative task transfer, continually learning new tasks without catastrophic forgetting, and improving compositional abilities.
-Wang et al.
-(2023b) conduct analysis on factual knowledge, reasoning, multilinguality, coding, and open-ended instruction following abilities of models trained with 12 instruction datasets and experimentally show that different instruction datasets may correspond to specific abilities.
-Also, it seems that winning across all evaluations using a single dataset or combination is challenging.
-3.4 Date-Efficient Learning
-Based on explorations of the impact of data quantity, data quality, and task composition on model performance discussed previously, some works propose to fine-tune LLM more efficiently with subset selection or specially designed fine-tuning strategy addressing different aspects of instruction data.
-Data Quantity
-AlShikh et al.
-(2023) introduce a metric to measure LLMs’ instruction-following ability.
-A binary classifier is used to predict whether the response is "answer-like" or not.
-Then, Instruction Following Score (IFS) is defined as the percentage of "answer-like" responses in the responses acquired by the whole instruction dataset.
-They suggest using this metric as an early stopping criterion without tuning on the full-sized dataset.
-Based on observations of different scaling patterns for different abilities, Dong et al.
-(2023) propose Dual-stage Mixed Fine-tuning (DMT) strategy to learn specialized abilities and general abilities sequentially while keeping a small proportion of specialized data to prevent forgetting.
-Data Quality
-Several works focus on selecting a subset of instruction data with the highest quality.
-Cao et al.
-(2023) adopt BlendSearch Wang et al.
-(2020) to automatically select the best subset.
-AlpaGasus Chen et al.
-(2023c) uses strong LLMs as auto-graders and selects data with scores above a threshold in the Alpaca dataset Taori et al.
-(2023).
-Motivated by the computational overhead caused by data pruning prior to fine-tuning, Attendu and Corbeil (2023) propose a dynamic data pruning method that periodically filters out unimportant examples during SFT using extended versions of EL2N metric Paul et al.
-(2021); Fayyaz et al.
-(2022).
-Without discarding data samples, OpenChat Wang et al.
-(2023a) considers the general SFT data as a mixture of a small amount of expert data and a large amount of sub-optimal data without any preference labels.
-Then, Conditioned-RLFT strategy is proposed, which treats different data sources as coarse-grained reward labels and optimizes the LLM as a class-conditioned policy.
-To enhance instruction diversity in the chosen subsets, DiverseEvol Wu et al.
-(2023) uses an iterative data sampling technique that selects new data points with maximized distances from any existing ones in model embedding space.
-Task Composition
-Given a small amount of target task data, Ivison et al.
-(2023) finds the relevant multitask subsets for fine-tuning according to the similarity between the pre-trained model’s representations.
-Similarly, Dynosaur Yin et al.
-(2023a) treats task selection based on instruction representations as a replay strategy in continual learning scenarios to mitigate catastrophic forgetting issues and improve generalization to unseen tasks.
-Yue et al.
-(2023) builds math generalist models MAmmoTH through instruction tuning on a unique hybrid of chain-of-thought and program-of-though rationales in math.
-Others
-Zhou et al.
-(2023b) introduce learnability as a new dimension of SFT data selection that data can be learned more effectively by the model are preferable and data lacking informative content or excessively demanding for the model should be avoided.
-They also propose LoBaSS to select SFT data using learnability as the principal criterion measured by the loss difference between fine-tuned and pre-trained models.
-Xu et al.
-(2023b) propose a contrastive post-training technique via contrastive pairs from LLMs with different levels of abilities.
-They also use a data curriculum scheme where the model learns progressively from the "easier part" to the "harder part".
-Data-Juicer Chen et al.
-(2023a) implements pipelines for LLM fine-tuning and operators for users to compose different data management recipes, as well as the evaluation of consequence model performance.
-4 Challenges and Future Directions
-The exploration of data management and its impact on LLM pretraining and SFT is still an ongoing task.
-In this section, we point out several challenges and corresponding future directions in training data management studies for LLMs.
-Comprehensive and Fine-grained Understanding
-As discussed in previous sections, many efforts have been made to understand the impacts of data management on different training stages addressing different aspects.
-While current studies contribute valuable pieces to the puzzle, a comprehensive understanding of the entire picture is still lacking.
-Moreover, explorations using different datasets and models on different tasks may lead to contradictory conclusions, e.g., the trade-off between quality and toxicity filtering Longpre et al.
-(2023b), detoxification v.s.
-debiasing Xu et al.
-(2021); Welbl et al.
-(2021), fine-tuning with a few high-quality data Zhou et al.
-(2023a); Chen et al.
-(2023b) v.s.
-data scaling Wei et al.
-(2021); Sanh et al.
-(2022), task composition Wang et al.
-(2022); Chung et al.
-(2022) v.s.
-expert models Jang et al.
-(2023); Wang et al.
-(2023b), etc.
-Hence, more fine-grained understanding is required to solve these conflicts.
-General Data Management Framework
-Although Data-Juicer Chen et al.
-(2023a) and Oasis Zhou et al.
-(2023c) propose data management systems to compose various data recipes in either the pretraining or SFT stage of LLM, practitioners still need to spend efforts on finding suitable datasets.
-Constructing a general data management framework suitable for a broad range of applications to reduce data management costs is an urgent and worthy future direction in developing and promoting LLMs.
-Data Efficiency Through Data Management
-Training more powerful LLMs with less and more effective training data is always an ongoing pursuit in the training of LLMs.
-As discussed in Section 3.4, works have been proposed to fine-tune LLMs to achieve various abilities in a data-efficient way.
-Further exploration in data management is expected to offer good opportunities for achieving better data efficiency.
-Data Curriculum
-Besides choosing better training data, data curriculum addressing the arrangement of data learning orders is also an important part of data management, e.g., learning from general abilities to target abilities or from easier tasks to harder tasks.
-There are a few works focusing on data curriculum in the training of LLMs Xu et al.
-(2023b); Dong et al.
-(2023); Yin et al.
-(2023a).
-Although effective in practice, there is still a lack of analysis of data curriculum strategies.
-Conflict Data Separation
-In the collection of training data, conflicts among the responses to the same query may exist.
-For example, given the same query, LLMs playing different roles may generate different responses.
-Mixing these samples together could lead to negative impacts on model performance because of the response conflicts.
-Thus, how to separate and effectively learn from these data samples is an interesting topic in the future.
-Multimodal Data Management
-Current research in data management mostly focuses on natural language processing.
-With the application of LLMs extending to multimodalities like vision, audio, and so on, the construction of multimodality datasets becomes more and more important.
-The proposed multi-modal LLMs usually construct their own instruction-tuning datasets collected from benchmark adaptation Zhang et al.
-(2023a); Gao et al.
-(2023) or self-instruction Pi et al.
-(2023); Yang et al.
-(2023b).
-The hybrid composition of language-only and multimodal data is also adopted in some works Dai et al.
-(2023); Zhao et al.
-(2023c).
-It is interesting to see the impacts of multimodal data management on the performance of fine-tuned multimodal LLMs, such as the data scaling law in multimodal instruction fine-tuning, the quality-control techniques in multimodal dataset construction, task balancing in multitask multimodal training, and so on.
-Hallucinations
+B.3 Hallucinations
 Despite their strong power, LLMs are notorious for their hallucinations, i.e.
@@
 Several works in hallucination trace down the occurrence of hallucination to the lack of pertinent knowledge and the internalization of false knowledge from the pretraining corpora Li et al.
-(2022); McKenna et al.
+(2022a); McKenna et al.
 (2023); Dziri et al.
@@
 It can be seen from the previous research that data management in both the pretraining and SFT stages can be a promising solution to hallucination.
-Social Biases and Fairness
-The problem of social biases and unfairness existing in current pretraining datasets and their impacts on pre-trained LLMs have been addressed in several works as discussed in Section 2.2.4.
-However, there is still a large gap between current prominent LLMs and ideal LLMs without social biases.
-Many questions are worth exploring, such as how to mitigate the potential biases in pretraining datasets, the existence of bias in the SFT datasets, and whether it is feasible to reduce social bias through SFT.
-5 Related Surveys
-As LLMs draw more and more attention, a handful of surveys have been published or preprinted addressing different aspects of the life circle of LLMs.
+Appendix C Related Surveys
+As LLMs draw more and more attention, a handful of surveys have been published or preprinted addressing different aspects of their development.
 Related to our work, several of them also include parts of the data preparation process in the pretraining or SFT of LLM.
@@
 (2023e) review the technologies aligning LLMs with human expectations including data collection, training methodologies, and model evaluation.
-Different from previous surveys, this survey provides a systematic and detailed overview of data management at both the pretraining and SFT stages of LLMs.
-We focus on the proper organization of training datasets and the effects of different data management strategies, providing a guiding resource for practitioners aiming to build powerful LLMs through effective and efficient data management.
-6 Conclusions
-This paper overviews data management in the training of LLMs.
-We discuss the pretraining and supervised fine-tuning stage of LLM successively and summarize the up-to-date research efforts into data quantity, data quality, and domain/task composition for each stage.
-Data management systems in the pretraining stage and Data-efficient learning in the supervised fine-tuning stage are also discussed.
-Finally, we highlight several challenges and promising future directions for LLM training data management.
-We hope this survey can provide insightful guidance for practitioners and inspire further research in effective and efficient data management for the development of LLMs.
+Unlike previous surveys, this survey provides a systematic and detailed overview of data management at both the pretraining and SFT stages of LLMs.
+We focus on the proper organization of training datasets and discuss recent research addressing the effects of different data management strategies, the evaluation of curated training datasets, and the latest advances in training data management strategies, providing a guiding resource for practitioners aiming to build powerful LLMs through efficient data management.
+Appendix D Comparison of Data Management Strategies Used by Representative LLMs
+We provide two summary tables, Table 1 for pretrained LLMs and Table 2 for SFT LLMs, with better and clearer comparison of the data management strategies used by current representative LLMs.
+Pretrained LLMs Open- sourced Quantity Deduplication Quality Filters Toxicity Filters Domian Composition
+T5  Raffel et al. (2020) $\surd$ 750GB N-gram Heuristic Heuristic 99% Web, < 1% Wiki
+GPT-3  Brown et al. (2020) 499B tokens MinHash, LSH Classifier 82% Web, 16% Books, 3% Wiki
+GLaM  Du et al. (2022) 1.6T tokens Classifier 46% Web, 28% Dialog, 20% Books, 6% Wiki
+LaMDA  Thoppilan et al. (2022) 1.56T words 50% Dialog, 25% Web, 12.5% Wiki, 12.5% Code
+Chinchilla  Hoffmann et al. (2022) 1.4T tokens N-gram, Doc-level Heuristic Heuristic 65% Web, 30% Books, 4% Code, 1% Wiki
+AlphaCode  Li et al. (2022b) 715.1GB Doc-level Heuristic 100% Code
+GLM  Zeng et al. (2022) $\surd$ 400B tokens 50% Pile, 50% Chinese Web data
+BLOOM  Le Scao et al. (2023) $\surd$ 1.61TB text SimHash, Substring clustering Heuristic Heuristic 60% Web, 10% Books, 10% Code, 10% Academic, 5% Dialog, 5% Wiki
+PaLM  Anil et al. (2023) 780B tokens Levenshtein distance Heuristic, Classifier Classifier 50% Dialog, 28% Web, 13% Books, 5% Code, 4% Wiki
+LLaMA  Touvron et al. (2023a) $\surd$ 1.4T tokens Line-level, Book-level Heuristic, Classifier Classifier 82% Web, 4.5% Code, 4.5% Wiki, 4.5% Books, 2.5% Academic, 2% Dialog
+Mistral  Jiang et al. (2023a) $\surd$ - - - - -
+phi-1/1.5  Gunasekar et al. (2023)  Li et al. (2023d) $\surd$ 7B tokens Classifier 99% Academic, <1% Code
+phi-2  Javaheripi and Bubeck (2023) $\surd$ 1.4B tokens Classifier
+GPT-4  OpenAI (2023) - - - - -
+LLaMA 2  Touvron et al. (2023b) $\surd$ 2.0T tokens Heuristic
+QWen  Bai et al. (2023) $\surd$ 3T tokens Exact Match, MinHash, LHS Heuristic, Classifier Classifier Web, Books, Codes, Academic
+Deepseek LLM  Bi et al. (2024) $\surd$ - - - - -
+Table 1: The data management strategies used by representative pretrained models.
+The blank units mean no specific design of corresponding strategies according to the original papers.
+The ’-’ means the data management process is not released.
+Part of the data is adopted from Longpre et al.
+(2023b)
+SFT LLMs Dataset Quantity Quality Control Diversity Control Complexity Enhancing No. of Tasks Task Balancing
+Tk-Instruct  Wang et al. (2022) NIv2 5M Heuristic Human 1616 Limited instances per task
+Flan-T5  Longpre et al. (2023a) Flan 2022 15M Input Inversion 1836 Experiments intuitions
+OPT-IML  Iyer et al. (2022) OPT-IML Bench 18M 2000 Experiments
+Alpaca  Taori et al. (2023) Alpaca 52K Heuristic ROUGE-L similarity 80
+Vicuna  Chiang et al. (2023) ShareGPT 70K Heuristic
+LIMA  Zhou et al. (2023a) LIMA 1K Heuristic Human Heuristic, Human
+Dolly  Conover et al. (2023) dolly-15k 15K Human
+Orca  Mukherjee et al. (2023) sampled Flan 2022 5M Chat-GPT/ GPT-4 augmentation
+WizardLM  Xu et al. (2023a) WizardCoder  Luo et al. (2023b) WizardMath  Luo et al. (2023a) WizardLM WizardCoder WizardMath 250K Evol-Instruct Evol-Instruct
+AlpaGasus  Chen et al. (2023c) AlpaGasus 9K Chat-GPT grading
+Platypus  Lee et al. (2023b) Open- Platypus 25K Dedup, Heuristic
+OpenChat  Wang et al. (2023a) ShareGPT 6K C-RLFT
+MAmmoTH  Yue et al. (2023) MathInstruct 260K 7 math fields Combining CoT and PoT
+Table 2: The data management strategies used by representative supervised finetuned models.
+The blank units mean no specific design of corresponding strategies according to the original papers.
+"NIv2" is the abbreviation for "Super-NaturalInstructions".
+"Dedup" is the abbreviation for "Deduplication".
+Appendix E Taxonomy
+The full taxonomy of research discussed in this survey is illustrated in Figure 3
+{forest}
+forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=center, font=, rectangle, draw=hidden-draw, rounded corners, align=left, text centered, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=13em,font=,, where level=2text width=14em,font=,, where level=3text width=11em,font=,, [ Data Management, ver [ Pretraining (§2), ver [ Domain Composition (§2.1) [ Longpre et al.
+(2023b), Nijkamp et al.
+(2023), Shen et al.
+(2023),
+Xie et al. (2023b), Xie et al. (2023a), Fan et al. (2023), Ye et al. (2024),
+Xia et al.
+(2023), leaf, text width=33em ] ] [ Data Quantity (§2.2) [ Scaling Laws [ Kaplan et al.
+(2020), Hoffmann et al.
+(2022), Su et al.
+(2024), leaf, text width=33em ] ] [ Data Repetition [ Villalobos et al.
+(2022), Muennighoff et al.
+(2023), Hernandez et al.
+(2022),
+Xue et al. (2023), Tirumala et al. (2023), leaf, text width=33em ] ] ] [ Data Quality (§2.3) [ Quality Filtering [ Gao (2021), Kreutzer et al. (2022), Gunasekar et al. (2023), Li et al. (2023d),
+Penedo et al. (2023), Marion et al. (2023), Longpre et al. (2023b),
+Kaddour (2023), Javaheripi and Bubeck (2023), Gan et al. (2023),
+Wettig et al. (2024), leaf, text width=33em ] ] [ Deduplication [ Lee et al. (2021), Kandpal et al. (2022), Silcock et al. (2022),
+Abbas et al. (2023), leaf, text width=33em ] ] [ Toxicity Filtering [ Luccioni and Viviano (2021), Xu et al. (2021), Welbl et al. (2021),
+Longpre et al.
+(2023b), leaf, text width=33em ] ] [ Diversity & Age [ Lee et al.
+(2023a), Maharana et al.
+(2023), Longpre et al.
+(2023b), leaf, text width=33em ] ] [ Social Bias* [ Dodge et al.
+(2021), Meade et al.
+(2022), Gururangan et al.
+(2022)
+Feng et al. (2023), leaf, text width=33em ] ] [ Hallucinations* [ Li et al. (2022a), McKenna et al. (2023), Dziri et al. (2022), leaf, text width=33em ] ] ] [ Relations Among
+Different Aspects (§2.4) [ Ge et al. (2024a), Goyal et al. (2024), Bi et al. (2024), Shen et al. (2023),
+Longpre et al. (2023b), leaf, text width=33em ] ] ] [ Supervised Fine-Tuning (§3), ver [ Task Composition (§3.1) [ Wei et al. (2021), Wang et al. (2022), Sanh et al. (2022), Chung et al. (2022),
+Longpre et al. (2023a), Jang et al. (2023), Chen et al. (2023b), Xia et al. (2024)
+Dong et al. (2023), Iyer et al. (2022), Wang et al. (2023b), Ivison et al. (2023),
+Lee et al. (2024), leaf, text width=33em ] ] [ Data Quality (§3.2) [ Instruction Quality [ Chia et al. (2023), Zhou et al. (2023a), Li et al. (2023b), Li et al. (2024b)
+Ding et al. (2023), Wang et al. (2023d), Li et al. (2023c), Zhou et al. (2023b),
+Cao et al. (2023), Madaan et al. (2023), Du et al. (2023), Li et al. (2024c)
+Lu et al. (2023a), Ye et al. (2023), Chen et al. (2023c), Li et al. (2023a),
+Li et al. (2023e), Bhatt et al. (2024), Chen and Mueller (2024),
+Yang et al. (2024), Mekala et al. (2024), He et al. (2024b), Liu et al. (2024), leaf, text width=33em ] ] [ Instruction Diversity [ Ding et al. (2023), Zhou et al. (2023a), Bukharin and Zhao (2023)
+Taori et al. (2023), Lu et al. (2023b), Wang et al. (2023c),
+Wan et al.
+(2023), Wu et al.
+(2023), Ge et al.
+(2024b), Huang et al.
+(2024), leaf, text width=33em ] ] [ Instruction Complexity [ Lu et al.
+(2023b), Xu et al.
+(2023a), Luo et al.
+(2023b), Mukherjee et al.
+(2023),
+Zhao et al. (2023b), He et al. (2023), Jiang et al. (2023b), Sun et al. (2024a)
+He et al. (2024a), leaf, text width=33em ] ] [ Prompt Design* [ Mishra et al. (2022), Khashabi et al. (2022), Gonen et al. (2022),
+Yin et al. (2023b), Kung and Peng (2023), Liang et al. (2023),
+Weber et al.
+(2023), Gudibande et al.
+(2023), Song et al.
+(2023), leaf, text width=33em ] ] [ Hallucinations* [ Zhou et al.
+(2023a), Chen et al.
+(2023c), Cao et al.
+(2023), Lee et al.
+(2023b), leaf, text width=33em ] ] ] [ Data Quantity (§3.3) [ Ji et al.
+(2023), Zhou et al.
+(2023a), Yuan et al.
+(2023), Zhang et al.
+(2024)
+Chen et al. (2023b), Dong et al. (2023), Song et al. (2023), leaf, text width=33em ] ] [ Dynamic
+Data-Efficient Learning (§3.4) [ Training Affects Data [ Attendu and Corbeil (2023), AlShikh et al.
+(2023), Kung et al.
+(2023), leaf, text width=33em ] ] [ Data Affects Training [ Yin et al.
+(2023a), Wang et al.
+(2023a), Dong et al.
+(2023),
+Xu et al. (2023b), Wang et al. (2024), Sun et al. (2024a), Kim and Lee (2024), leaf, text width=33em ] ] ] [ Relations Among
+Different Aspects (§2.4) [ Lu et al. (2023b), Bukharin and Zhao (2023), Xu et al. (2023c), leaf, text width=33em ] ] ] ]
+Figure 3: Taxonomy of research in data management for pretraining and supervised fine-tuning of Large Language Models (LLM).